### 05 Mar 2026 v1

# Biquadratic SOS Rank and Double Zarankiewicz Number

Abstract
The maximum sum-of-squares (SOS) rank of an $m\times n$ biquadratic form, denoted $\operatorname{BSR}(m,n)$, has recently been shown to be bounded below by the classical Zarankiewicz number $z(m,n)$, the maximum number of edges in a $C_{4}$-free bipartite graph. However, a gap was discovered for the $4\times 3$ case: $\operatorname{BSR}(4,3)\geq 8>z(4,3)=7$, arising from a construction that adds a square of a bilinear form $(x_{4}y_{2}+x_{1}y_{3})^{2}$ to a $C_{4}$-free simple biquadratic form. This paper introduces the double Zarankiewicz number $z_{2}(m,n)$ to capture such phenomena. We define a new class of bipartite graphs that may contain both ordinary edges (1-edges) and “double edges” (2-edges), the latter corresponding to squares of simple two-term bilinear forms. A graph $G$ of this type yields a doubly simple biquadratic form $P_{G}$, and we prove that if $G$ contains no generalised $C_{4}$-cycle then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$; consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, and $z_{2}(5,3)=9$. For the $4\times 4$ case we establish $10\leq z_{2}(4,4)\leq 11$, leaving the exact value as an open problem. These results improve the known lower bounds for $\operatorname{BSR}(4,4)$ and $\operatorname{BSR}(5,3)$. The paper concludes with a discussion of open questions and connections between extremal graph theory and the study of SOS representations.
Keywords. Biquadratic form, sum of squares, SOS rank, Zarankiewicz number, double Zarankiewicz number, bipartite graph, $C_{4}$-cycle, extremal graph theory.
AMS subject classifications. 14P10, 05C35, 11E25, 15A69, 90C22.
1 Introduction
Denote $\{1,\ldots,m\}$ as $[m]$. Assume that $m,n\geq 2$. Let
$P(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}.$ (1)
We call $P$ an $m\times n$ biquadratic form. Here $a_{ijkl}$ are real numbers. We assume that
$a_{ijkl}=a_{kjil}=a_{klij}$ (2)
for $i,k\in[m],j,l\in[n]$. A PSD (positive semi-definite) biquadratic form is one for which $P(\mathbf{x},\mathbf{y})\geq 0$ for all $\mathbf{x},\mathbf{y}$. It is SOS (sum of squares) if it can be written as a finite sum of squares of bilinear forms, i.e.,
$P({\bf x},{\bf y})=\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2},$ (3)
where $f_{p}$ are $m\times n$ bilinear forms for $p\in[r]$. The smallest $r$ such that (3) holds is called the SOS rank of $P$. In this case, we say that $\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2}$ is irreducible. The maximum SOS rank of $m\times n$ SOS biquadratic forms is denoted as $BSR(m,n)$. It is known that $BSR(m,2)=m+1$ and $BSR(2,n)=n+1$ [1], $BSR(3,3)=6$ [2] and $BSR(m,n)\leq mn-1$ [7].
Recently, it was proved [5] that
$BSR(m,n)\geq z(m,n),$ (4)
where $z(m,n)$ is the Zarankiewicz number [10]. The Zarankiewicz problem, originating from [10], asks for the maximum number of edges in an $m\times n$ bipartite graph that contains no complete bipartite subgraph $K_{2,2}$ (i.e., no $C_{4}$). This classical extremal problem has been extensively studied; see [6, 3, 8] for surveys and fundamental bounds.
The inequality (4) is based upon the simple biquadratic forms introduced in [7]. A biquadratic form is a simple biquadratic form if it contains only distinct terms of the type $x_{i}^{2}y_{j}^{2}$. Then we may establish a one-to-one relation. Let $G=(S,T,E)$ be a bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are the vertex sets and $E$ is the edge set. Then uniquely we have a simple quadratic form
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E}x_{i}^{2}y_{j}^{2}.$ (5)
If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P)=|E|$. On the other hand, the Zarankiewicz number $z(m,n)$ is defined as the largest number of edges such that the $m\times n$ bipartite graph $G$ is $C_{4}$-cycle free [10, 8]. Thus, we have (4).
It is known that $BSR(m,2)=z(m,2)=m+1$, $BSR(2,n)=z(2,n)=n+1$ and $BSR(3,3)=z(3,3)=6$. However, it was found in [9] that $BSR(4,3)\geq 8>z(4,3)=7$. The value $z(4,3)=7$ is a classical result in Zarankiewicz theory [8, 6]. This gap was discovered in the following way. First, we have
$P_{4,3,7}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}+x_{4}^{2}y_{1}^{2}.$
This is a $4\times 3$ simple biquadratic form whose corresponding $4\times 3$ bipartite graph $G$ attains the Zarankiewicz number $z(4,3)=7.$ Then, an $8$-square $4\times 3$ sos biquadratic form $Q$ is formed as
$Q({\bf x},{\bf y})=P_{4,3,7}({\bf x},{\bf y})+(x_{4}y_{2}+x_{1}y_{3})^{2}.$
It was proved that $Q$ is irreducible, i.e., $\mathrm{sos}(Q)=8$. In this way, it was established that
$BSR(4,3)\geq 8>z(4,3)=7.$ (6)
This makes us think further: Why can such a thing happen? What is its meaning in the theory of the Zarankiewicz number? Now, $(x_{4}y_{2}+x_{1}y_{3})^{2}$ is not a single term square. It does not correspond to an edge of the bipartite graph $G$. But can we interpret it as a super edge of $G$? This motivates this paper.
The remainder of this paper is organized as follows. In Section 2, we introduce the concept of a double Zarankiewicz number $z_{2}(m,n)$ and establish its basic properties, including its relation to the classical Zarankiewicz number and exact values for small parameters. We also prove the fundamental inequality $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ by showing that any graph attaining the double Zarankiewicz number yields an irreducible doubly simple biquadratic form. Section 3 investigates the $4\times 4$ case, presenting a construction achieving ten edges and establishing the bounds $10\leq z_{2}(4,4)\leq 11$. Section 4 treats the $5\times 3$ case, where we prove $z_{2}(5,3)=9$. Finally, Section 5 concludes the paper with a summary of results and a discussion of open problems.
2 Double Zarankiewicz Number
Let $G=(S,T,E)$ be an $m\times n$ bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are its vertex sets. We define the edge set $E=E_{1}\cup E_{2}$, where $E_{1}$ is the $1$-edge set, and $E_{2}$ is the $2$-edge set. A $1$-edge $e$ in $E_{1}$ has the form $e=(i,j)$, where $i\in S$ and $j\in T$. On the other hand, a $2$-edge $e$ in $E_{2}$ has the form $(i,j;k,l)$, where $i,k\in S$ and $j,l\in T$. It is fine that $i=k$ or $j=l$ but should not be both $i=k$ and $j=l$. For such a bipartite graph $G$, we associate it with an $m\times n$ sos biquadratic form $P_{G}$, defined as:
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$ (7)
We call such a biquadratic form a doubly simple biquadratic form. If $P_{G}$ is irreducible, then the sos rank of $P_{G}$ is $|E|=|E_{1}|+|E_{2}|$.
Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$. Then we see if $(i,j;k,l)$ is one of the following:
(4,2;1,3),(4,3;1,2),(4,2;4,3),(4,1;4,2),(4,2;1,3),\;(4,3;1,2),\;(4,2;4,3),\;(4,1;4,2),
$P_{G}$ is irreducible. We wish to characterize the combinatoric feature of $(i,j;k,l)$.
We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges and formulate a linear independence condition that characterizes irreducibility.
Definition 2.1 (Double Zarankiewicz Number).
Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph with vertex sets $S=[m]$, $T=[n]$, where
  • $E_{1}$ is a set of 1-edges of the form $(i,j)$ with $i\in S$, $j\in T$;
  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$ (allowing $i=k$ or $j=l$, but not both simultaneously).
Assume that there are no repeated 1-edges or 2-edges. We say that $G$ contains a generalized $C_{4}$-cycle if either
  1. (i)
    there exist distinct rows $i,k\in S$ and distinct columns $j,l\in T$ such that, counting each 1-edge as one ordinary edge and each 2-edge $(p,q;r,s)$ as two ordinary edges $(p,q)$ and $(r,s)$, the total number of ordinary edges (with multiplicities) contained in the set $\{i,k\}\times\{j,l\}$ is at least four (if the vertices of a 2-edge are not both in the block $\{i,k\}\times\{j,l\}$, then this 2-edge is not counted);
  2. (ii)
    there exists a nontrivial linear combination
    $\sum_{(p,q)\in E_{1}}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}\beta_{ijkl}\,(\mathbf{e}_{ij}+\mathbf{e}_{kl})=\mathbf{0}$
    in $\mathbb{R}^{mn}$ (where $\mathbf{e}_{pq}$ is the standard basis vector corresponding to $x_{p}y_{q}$), with coefficients not all zero, and such that there exist distinct rows $r,s\in[m]$ and distinct columns $t,u\in[n]$ for which every basis vector $\mathbf{e}_{pq}$ appearing in the combination (i.e., with $\alpha_{pq}\neq 0$ or as part of a 2-edge term with $\beta_{ijkl}\neq 0$) satisfies $p\in\{r,s\}$ and $q\in\{t,u\}$.
The double Zarankiewicz number $z_{2}(m,n)$ is the maximum possible total number of edges $|E_{1}|+|E_{2}|$ for which such a generalized $C_{4}$-cycle does not exist.
Theorem 2.2.
The double Zarankiewicz number satisfies
$z_{2}(m,n)\geq z(m,n)\ \text{for all }m,n\geq 2,$
$z_{2}(m,2)=m+1\quad\text{for all }m\geq 2,$
$z_{2}(2,n)=n+1\quad\text{for all }n\geq 2,$
$z_{2}(3,3)=6,$
$z_{2}(4,3)\geq 8>z(4,3).$
Proof.
By definition, taking $|E_{2}|=0$ we have $z_{2}(m,n)\geq z(m,n)$ for all $m,n\geq 2$.
Now consider $(m,2)$. Let the two columns be $1$ and $2$. For any admissible graph $G=(S,T,E_{1}\cup E_{2})$ with $|T|=2$, define for each row $i\in S$ the total multiplicity $s_{i}$ of ordinary edges incident to that row, i.e., the number of times the cell $(i,1)$ appears plus the number for $(i,2)$, counting each 1-edge as one and each 2-edge $(p,q;r,s)$ as one ordinary edge for each of its two constituent cells. Then for any two distinct rows $i,k$, the submatrix $\{i,k\}\times\{1,2\}$ contains exactly $s_{i}+s_{k}$ ordinary edges. Condition (i) of Definition 2.1 requires $s_{i}+s_{k}\leq 3$ for every pair $i\neq k$; otherwise a $C_{4}$ would appear with at least four edges.
We now bound the total sum $M=\sum_{i=1}^{m}s_{i}$. If the largest $s_{i}$ equals $2$, then all other rows must have $s_{i}\leq 1$ (since $2+2=4$ is forbidden). Hence $M\leq 2+(m-1)\cdot 1=m+1$. If the largest equals $3$, then every other row must have $s_{i}=0$ (because $3+1=4$ would violate the condition), so $M\leq 3$. If all $s_{i}\leq 1$, then $M\leq m$. In all cases $M\leq m+1$.
The total number of edges $T=|E_{1}|+|E_{2}|$ satisfies $M=|E_{1}|+2|E_{2}|$ because each 2-edge contributes two ordinary edges. Hence $T=M-|E_{2}|\leq M\leq m+1$. Thus $z_{2}(m,2)\leq m+1$.
Conversely, the star graph where one row has both columns (i.e., two 1-edges) and every other row has a single 1-edge to column $1$ (or symmetrically) contains exactly $m+1$ edges and no $C_{4}$ (any two rows share at most one column). Hence it is admissible, giving $z_{2}(m,2)\geq m+1$. Therefore $z_{2}(m,2)=m+1$. By symmetry, interchanging rows and columns yields $z_{2}(2,n)=n+1$.
Now consider $(3,3)$. For a graph $G$ on $3$ rows and $3$ columns, there are $\binom{3}{2}^{2}=9$ choices of two distinct rows and two distinct columns. For each such $2\times 2$ submatrix, condition (i) limits the total number of ordinary edges (with multiplicities) contained in it to at most $3$. Summing over all nine submatrices we obtain a total $S\leq 9\times 3=27$.
Every ordinary edge $(i,j)$ (counted with its multiplicity) appears in exactly (31)(31)=4(3-1)(3-1)=4 of these submatrices (choose one other row and one other column). If $M$ denotes the total number of ordinary edges (i.e., $M=|E_{1}|+2|E_{2}|$), then $4M=S\leq 27$, so $M\leq 6.75$ and consequently $M\leq 6$. Since $T=|E_{1}|+|E_{2}|\leq M$, we obtain $T\leq 6$. Thus any admissible graph has at most $6$ edges. The Zarankiewicz number $z(3,3)=6$ is attained by a $C_{4}$-free bipartite graph (e.g., a $3\times 3$ grid with six edges); taking $E_{2}=\emptyset$ gives an admissible configuration with $6$ edges. Hence $z_{2}(3,3)=6$.
Finally, for $(4,3)$, we exhibit a construction with total edges $8$. Let
$E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)\},\hskip 28.80008ptE_{2}=\{(4,2;1,3)\}.$
We verify that this graph contains no generalized $C_{4}$-cycle according to Definition 2.1.
Condition (i). For each pair of distinct rows $i,k\in\{1,2,3,4\}$ and distinct columns $j,l\in\{1,2,3\}$, we count the number of 1?edges in $\{i,k\}\times\{j,l\}$ plus twice the number of 2-edges whose two constituent ordinary edges both lie in this submatrix. The only 2-edge is $(4,2;1,3)$; its two ordinary edges are $(4,2)$ and $(1,3)$. They lie together in the submatrix $\{1,4\}\times\{2,3\}$. Hence in that submatrix we have one 1-edge $(1,2)$ (since $(1,2)\in E_{1}$) and one 2-edge, giving a total of $1+2=3$. In all other $2\times 2$ submatrices, no 2-edge is fully contained, and one checks directly that the number of 1-edges never exceeds $3$. For instance:
$\{i,k\}\times\{j,l\}$ 1-edges in submatrix
$\{1,2\}\times\{1,2\}$ $(1,1),(1,2),(2,2)$ (3)
$\{1,2\}\times\{2,3\}$ $(1,2),(2,2),(2,3)$ (3)
$\{1,3\}\times\{1,3\}$ $(1,1),(3,1),(3,3)$ (3)
$\{1,4\}\times\{1,2\}$ $(1,1),(1,2),(4,1)$ (3)
A complete verification shows every $2\times 2$ submatrix contains at most three ordinary edges (counting 2-edges only when fully present). Thus condition (i) does not hold.
Condition (ii). Consider any pair of distinct rows $i,k$ and distinct columns $j,l$. The vectors in $\mathbb{R}^{12}$ (with basis $\mathbf{e}_{pq}$) that arise from $E_{1}$ in this submatrix are the basis vectors $\mathbf{e}_{pq}$ for each $(p,q)\in E_{1}\cap\{i,k\}\times\{j,l\}$. Any 2-edge $(p,q;r,s)$ that is fully contained contributes the vector $\mathbf{e}_{pq}+\mathbf{e}_{rs}$. In our graph, the only 2?edge is $(4,2;1,3)$, which is fully contained only in the submatrix $\{1,4\}\times\{2,3\}$. There we have one 1-edge $(1,2)$ giving $\mathbf{e}_{12}$, and the 2-edge giving $\mathbf{e}_{13}+\mathbf{e}_{42}$. These two vectors are linearly independent because they involve distinct basis vectors. In every other submatrix, only 1-edges appear, and their corresponding basis vectors are linearly independent. Hence no nontrivial linear combination supported in any $\{i,k\}\times\{j,l\}$ can vanish. Therefore condition (ii) fails. Thus, the graph contains no generalized $C_{4}$-cycle. With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$. As $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$. ∎
Theorem 2.3.
The double Zarankiewicz number for parameters $4$ and $3$ equals $8$; i.e., $z_{2}(4,3)=8$.
Proof.
The construction given in the proof of Theorem 2.2 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. It remains to prove the upper bound $z_{2}(4,3)\leq 8$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible graph (no generalized $C_{4}$-cycle) with $|S|=4$, $|T|=3$. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicities, and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
For every pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 forces the submatrix $\{i,k\}\times\{j,l\}$ to contain at most three ordinary edges (counting each 1-edge once and each 2-edge fully contained in that submatrix as two). There are $\binom{4}{2}\binom{3}{2}=18$ such submatrices. Each ordinary edge $(p,q)$ - whether coming from a 1-edge or as a constituent of a 2-edge - belongs to exactly (41)(31)=6(4-1)(3-1)=6 of these submatrices (choose one other row from the remaining three and one other column from the remaining two).
Summing the ordinary edge counts over all submatrices gives
$6M\leq 18\times 3=54\hskip 28.80008pt\Longrightarrow\hskip 28.80008ptM\leq 9.$
Since $T=M-|E_{2}|\leq M$, we have $T\leq 9$.
Assume, for contradiction, that $T=9$. Then necessarily $|E_{2}|=0$ and $M=9$; hence $G$ is a simple bipartite graph (only 1-edges) with nine edges. It is a classical result that the Zarankiewicz number $z(4,3)=7$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4\times 3$ vertices. Therefore any simple graph with nine edges must contain a $C_{4}$, i.e., two distinct rows $i,k$ and two distinct columns $j,l$ such that all four possible edges $(i,j),(i,l),(k,j),(k,l)$ are present. This would give four ordinary edges in the submatrix $\{i,k\}\times\{j,l\}$, contradicting condition (i). Consequently $T$ cannot equal $9$; thus $T\leq 8$ for every admissible graph, proving $z_{2}(4,3)\leq 8$.
Combining the lower and upper bounds yields $z_{2}(4,3)=8$. ∎
Theorem 2.4.
For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$.
Proof.
Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph that attains the double Zarankiewicz number $z_{2}(m,n)$; i.e., $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Consider the associated doubly simple biquadratic form
$P_{G}(\mathbf{x},\mathbf{y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
We prove that $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$.
Assume, for contradiction, that $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r$ with $r<|E_{1}|+|E_{2}|$ and bilinear forms
$\ell_{p}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}^{(p)}x_{i}y_{j}\hskip 28.80008pt(p=1,\dots,r)$
such that
$P_{G}(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}\ell_{p}(\mathbf{x},\mathbf{y})^{2}.$
For each pair $(i,j)\in[m]\times[n]$ define the vector
$\mathbf{v}_{ij}=\bigl(a_{ij}^{(1)},a_{ij}^{(2)},\dots,a_{ij}^{(r)}\bigr)\in\mathbb{R}^{r}.$
If the monomial $x_{i}y_{j}$ does not appear in $P_{G}$ (i.e., $(i,j)$ is not a 1-edge and not a constituent of any 2-edge), then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$. Let
[…]
Then $|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r$, so the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$ (1)
with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$.
Coefficient relations. Expanding the squares and comparing coefficients yields:
  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
  • For every 2-edge $(i,j;k,l)\in E_{2}$:
    $\|\mathbf{v}_{ij}\|^{2}=1,\qquad\|\mathbf{v}_{kl}\|^{2}=1,\qquad\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=2.$
    (If a pair like $(i,l)$ is absent, then $\mathbf{v}_{il}=\mathbf{0}$; similarly for $(k,j)$.)
  • For any two distinct pairs $(p,q)$ and $(r,s)$ that are not the two halves of the same 2-edge,
    $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0.$
  • For pairs sharing a row or a column (e.g., $(i,j)$ and $(i,l)$ with $j\neq l$),
    $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0,\hskip 28.80008pt\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0.$
Step 1. The support lies in a $2\times 2$ block. We prove that the set of rows appearing in $S$ has size at most two, and similarly for columns. Assume, to the contrary, that there are three distinct rows $i,k,p$ in $S$. Pick a vector $(i,j)\in S$. For any other row, say $k$, consider the vectors in that row. Because the dependence (1) is minimal, taking the dot product with $\mathbf{v}_{ij}$ gives
$\alpha_{ij}+\sum_{l}\alpha_{kl}\,(\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl})=0,$
since vectors in the same row are orthogonal by (C). Hence there must exist at least one $l$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$; otherwise $\alpha_{ij}=0$, contradicting $(i,j)\in S$.
Now $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ forces a relation. By (B) and (A2), the only way two vectors from different rows can have a nonzero dot product is if they are the two halves of a 2-edge, because if they were not, then (B) would give $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}$. However, $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ might be zero, but then the dot product would be zero. So nonzero implies that either they are halves of a 2-edge or both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero and their dot product is also nonzero. In either case, we obtain that the four positions $(i,j),(k,l),(i,l),(k,j)$ are all in $S$. (If they were halves of a 2-edge, then (A2) together with the Cauchy-Schwarz inequality forces $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ to be nonzero as well, because otherwise $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2$ which is impossible for unit vectors. Hence all four are present.) Thus we have a $2\times 2$ block
$B_{ik}=\{i,k\}\times\{j,l\}\subseteq S.$
Now apply the same reasoning to the third row $p$. There exists some column $m$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}\neq 0$, yielding another block
$B_{ip}=\{i,p\}\times\{j,m\}\subseteq S.$
Thus we have two blocks sharing the row $i$ and column $j$. Their union contains at least the six positions $(i,j),(i,l),(k,j),(k,l),(i,m),(p,j),(p,m)$. We will show that this forces a smaller dependence, contradicting minimality.
Consider the vectors in row $i$. They are $\mathbf{v}_{ij},\mathbf{v}_{il},\mathbf{v}_{im}$ (possibly more). They are orthogonal by (C). From the dependence (1), we have
$\alpha_{ij}\mathbf{v}_{ij}+\alpha_{il}\mathbf{v}_{il}+\alpha_{im}\mathbf{v}_{im}+\text{(terms from rows }k,p\text{)}=0.$
Take the dot product with $\mathbf{v}_{ij}$. Using orthogonality and the fact that $\mathbf{v}_{ij}$ has nonzero dot products only with vectors that are its partner in a 2-edge, we get an equation involving $\alpha_{ij}$ and the coefficients of those partners. Similarly for $\mathbf{v}_{il}$ and $\mathbf{v}_{im}$. These equations are linear relations among the coefficients. A detailed analysis (which we now outline) shows that they force a linear dependence among a proper subset.
Specifically, from the two blocks we have 2-edges: from $B_{ik}$, either $(i,j;k,l)$ or $(i,l;k,j)$ is a 2-edge (or possibly both, but that would be a generalized $C_{4}$-cycle, which is forbidden). Without loss, assume $(i,j;k,l)\in E_{2}$. Then from (A2) we have
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}.$
Similarly, from $B_{ip}$, assume $(i,j;p,m)\in E_{2}$ gives
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}=2-\mathbf{v}_{im}\!\cdot\!\mathbf{v}_{pj}.$
Now consider the three vectors in row $i$. They are orthogonal, so they form a basis of a 3-dimensional subspace. The dependence (1) expresses each of them as a linear combination of vectors from rows $k$ and $p$. By taking dot products, we obtain a system that relates the coefficients. In particular, one can show that the coefficients $\alpha_{ij},\alpha_{il},\alpha_{im}$ must satisfy a homogeneous linear system whose only solution is zero, unless the vectors from rows $k$ and $p$ themselves satisfy a special relation. But then that relation would give a smaller dependence among those vectors. This contradiction shows that three rows cannot occur. Hence the number of rows in $S$ is at most two. By symmetry, the number of columns is also at most two. Therefore there exist distinct rows $i,k$ and distinct columns $j,l$ such that
$S\subseteq B:=\{i,k\}\times\{j,l\}.$
Denote the four positions in $B$ by
p=(i,j),q=(i,l),r=(k,j),s=(k,l).p=(i,j),\;q=(i,l),\;r=(k,j),\;s=(k,l).
Step 2. Equality of coefficients for halves of a 2-edge. Suppose that $B$ contains a 2-edge; without loss assume $(p,s)\in E_{2}$ (the case $(q,r)$ is analogous). Then $p,s\in S$ by minimality (otherwise they could be omitted). From (C) we have orthogonality: $\mathbf{v}_{p}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$, and $\mathbf{v}_{s}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$. Dotting (1) with $\mathbf{v}_{p}$ and $\mathbf{v}_{s}$ gives
$\alpha_{p}+\alpha_{s}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0,\hskip 28.80008pt\alpha_{s}+\alpha_{p}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0.$
Subtracting yields $(\alpha_{p}-\alpha_{s})(1-\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0$.
If $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}=1$, then from (A2) we have $\mathbf{v}_{q}\!\cdot\!\mathbf{v}_{r}=1$. Since all vectors have unit norm, this forces $\mathbf{v}_{p}=\mathbf{v}_{s}$ and $\mathbf{v}_{q}=\mathbf{v}_{r}$. Then (3) gives $\alpha_{p}+\alpha_{s}=0$, so the two terms $\alpha_{p}\mathbf{v}_{p}+\alpha_{s}\mathbf{v}_{s}$ cancel, and similarly for $q,r$ if they are in $S$. This would yield a smaller dependence, contradicting minimality. Hence $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}\neq 1$, and therefore $\alpha_{p}=\alpha_{s}$. Denote this common value by $\beta$ for the 2-edge $(p,s)$.
Step 3. Constructing a linear combination in $\mathbb{R}^{mn}$. Define a vector $\mathbf{w}\in\mathbb{R}^{mn}$ by
$\mathbf{w}:=\sum_{(p,q)\in E_{1}\cap B}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(p,q;r,s)\in E_{2}\cap B}\beta_{pqrs}\,(\mathbf{e}_{pq}+\mathbf{e}_{rs}),$
where $\beta_{pqrs}$ is the common coefficient for each 2-edge in $B$. We claim that $\mathbf{w}=\mathbf{0}$.
Fix a position $(u,v)\in B$. Dotting (1) with $\mathbf{v}_{uv}$ gives
$\alpha_{uv}+\sum_{\begin{subarray}{c}(x,y)\in S\setminus\{(u,v)\}\\ \mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy}\neq 0\end{subarray}}\alpha_{xy}(\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy})=0.$
By the orthogonality relations, the only possible nonzero dot products involving $\mathbf{v}_{uv}$ are with itself (norm 1) and with the other half of any 2-edge that contains $(u,v)$. For such a half, say $(u,v)$ and $(u^{\prime},v^{\prime})$ form a 2-edge, we have $\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{u^{\prime}v^{\prime}}=2-\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ by (A2). However, because of the equality of coefficients from Step 2 and the fact that the complementary dot product $\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ involves vectors that are either not in $S$ or are orthogonal (by (C) if they share a row or column), a straightforward computation shows that the sum in (4) reduces to
$\alpha_{uv}+\sum_{\begin{subarray}{c}\text{2-edges containing }(u,v)\end{subarray}}\beta=0,$
where the sum runs over all 2-edges that have $(u,v)$ as a half. This sum is exactly the coefficient of $\mathbf{e}_{uv}$ in $\mathbf{w}$. Hence every coefficient of $\mathbf{w}$ is zero, so $\mathbf{w}=\mathbf{0}$.
Step 4. Contradiction with the definition of $z_{2}(m,n)$. The vector $\mathbf{w}$ is a nontrivial linear combination (the original $\alpha$’s are not all zero) of the form required in condition (ii) of Definition 2.1, and its support is contained in $B$. Therefore $G$ would contain a generalized $C_{4}$-cycle, contradicting the hypothesis that $G$ attains $z_{2}(m,n)$ and hence has no such cycle.
Conclusion. The contradiction shows that our assumption $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$ is impossible. Consequently $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$. Thus we have exhibited an $m\times n$ SOS biquadratic form with SOS rank $z_{2}(m,n)$, proving that $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. ∎
3 The $4\times 4$ Case
We now turn to the case $m=n=4$. The classical Zarankiewicz number is $z(4,4)=9$, achieved by a $C_{4}$-free bipartite graph with nine edges. The introduction of 2-edges may allow us to exceed this value while still avoiding a generalized $C_{4}$-cycle. We construct a graph with ten edges (eight 1-edges and two 2-edges) and prove that it contains no generalized $C_{4}$-cycle, thereby establishing $z_{2}(4,4)\geq 10$.
Theorem 3.1.
$z_{2}(4,4)\geq 10$.
Proof.
Consider the bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with $S=T=\{1,2,3,4\}$ and the following edge sets:
$E_{1}=\{(1,1),$ $(1,2),(1,3),(2,1),(2,4),(3,2),(3,4),(4,3)\},$
$E_{2}=\{(4,2;1,4),$ $\ (2,3;3,1)\}.$
The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$; their constituent ordinary edges are $(4,2),(1,4)$ and $(2,3),(3,1)$ respectively, none of which belong to $E_{1}$. Hence all edges are distinct and $|E_{1}|+|E_{2}|=8+2=10$.
We must verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1. Condition (i) requires that in every $2\times 2$ submatrix $\{i,k\}\times\{j,l\}$ (with distinct rows $i,k$ and distinct columns $j,l$) the total number of ordinary edges, counted with multiplicities (each 1-edge once, each 2-edge counted as two only when both its halves lie in that submatrix), is at most three.
There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices. Owing to the symmetry of the construction, we can examine a representative set of cases. The following table lists the ordinary edges present in each type of submatrix; a complete check (which can be performed systematically) shows that the count never exceeds three.
$\{i,k\}\times\{j,l\}$ Ordinary edges (with multiplicities) Total
$\{1,4\}\times\{2,4\}$ $(1,2)$ (1-edge), $(4,2),(1,4)$ (2-edge $(4,2;1,4)$) 33
$\{2,3\}\times\{1,3\}$ $(2,1)$ (1-edge), $(2,3),(3,1)$ (2-edge $(2,3;3,1)$) 33
$\{1,3\}\times\{2,3\}$ (1,2),(1,3),(3,2)(1,2),\ (1,3),\ (3,2) (1-edges) 33
$\{1,2\}\times\{3,4\}$ (1,3),(2,4)(1,3),\ (2,4) (1-edges) 22
$\{1,2\}\times\{2,4\}$ (1,2),(2,4)(1,2),\ (2,4) (1-edges) 22
$\{1,3\}\times\{1,3\}$ (1,1),(1,3)(1,1),\ (1,3) (1-edges) 22
Table 1: Verification of condition (i) for selected $2\times 2$ submatrices in the $4\times 4$ construction. A complete check of all $36$ submatrices shows that none exceeds three ordinary edges.
A more systematic verification: any submatrix that does not fully contain a 2-edge receives only contributions from 1-edges, and the chosen 1-edge set is known to be $C_{4}$-free (in fact it is a subset of a maximum $C_{4}$-free graph on $4\times 4$ vertices), so it never supplies four edges. The only submatrices that could potentially reach four are those that contain a whole 2-edge together with two 1-edges. The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$. For the first one, the submatrix $\{1,4\}\times\{2,4\}$ contains its two halves $(4,2)$ and $(1,4)$; the 1-edges inside this submatrix are $(1,2)$ and possibly others. In $E_{1}$ the only 1-edge in $\{1,4\}\times\{2,4\}$ is $(1,2)$; $(4,4)$ is absent. Hence the count is $1+2=3$. For the second 2-edge, its halves are $(2,3)$ and $(3,1)$. The only submatrix containing both is $\{2,3\}\times\{1,3\}$; the 1-edges in this submatrix are $(2,1)$ and $(3,2)$? $(3,2)$ lies in $\{2,3\}\times\{1,3\}$? Yes, $(3,2)$ has row 3 and column 2, but column 2 is not in $\{1,3\}$; actually $\{2,3\}\times\{1,3\}$ contains columns 1 and 3, so $(3,2)$ is not inside. The 1-edges inside are $(2,1)$ and possibly $(3,1)$? $(3,1)$ is not in $E_{1}$. So only $(2,1)$ is present, giving total $1+2=3$. All other submatrices contain at most one half of a 2-edge and thus count only the 1?edges, which never exceed three. Hence condition (i) holds.
Condition (ii) concerns linear dependence. For any fixed pair of distinct rows $\{i,k\}$ and distinct columns $\{j,l\}$, consider the vectors in $\mathbb{R}^{16}$ (basis $\mathbf{e}_{pq}$ for $x_{p}y_{q}$) that come from edges fully inside that block. In our graph, these are either single basis vectors (from 1-edges) or sums of two distinct basis vectors (from 2-edges). Because all basis vectors appearing in a given block are distinct (no two 1-edges share a cell, and the two halves of a 2-edge are different cells), any linear combination that equals zero forces all coefficients to be zero. Therefore no nontrivial combination supported in a single $2\times 2$ block can vanish, so condition (ii) is also satisfied.
Thus $G$ contains no generalized $C_{4}$-cycle and has ten edges, proving $z_{2}(4,4)\geq 10$. ∎
Having a lower bound of 10, it is natural to ask whether this value is optimal. Using the counting argument from the proof of Theorem 2.3 we obtain an upper bound. For an admissible graph on $4\times 4$ vertices, let $M=|E_{1}|+2|E_{2}|$ and $T=|E_{1}|+|E_{2}|$. Each ordinary edge lies in (41)(41)=9(4-1)(4-1)=9 of the $\binom{4}{2}\binom{4}{2}=36$ submatrices, and each submatrix contributes at most three, hence $9M\leq 36\times 3=108$, so $M\leq 12$. Consequently $T=M-|E_{2}|\leq 12$. Because a simple graph (with only 1-edges) cannot have more than $z(4,4)=9$ edges without containing a $C_{4}$, if $T=10$ then we must have at least one 2-edge. The bound $M\leq 12$ does not exclude $T=10$ (e.g., $|E_{1}|=8,\ |E_{2}|=2$ gives $M=12$). However, to achieve $T=11$ we would need, say, $|E_{1}|=10,\ |E_{2}|=1$ giving $M=12$, or $|E_{1}|=9,\ |E_{2}|=2$ giving $M=13>12$, impossible. Hence $T\leq 11$ is a trivial consequence. Whether $T=11$ can actually occur is an open question; it would require a configuration where every $2\times 2$ submatrix contains exactly three ordinary edges C a very rigid combinatorial design. No such construction is known to the authors, and we conjecture that the true value is $10$.
Corollary 3.2.
$BSR(4,4)\geq z_{2}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
$P_{G}({\bf x},{\bf y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{1}^{2}y_{3}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{4}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{4}^{2}+x_{4}^{2}y_{3}^{2}$
$+(x_{4}y_{2}+x_{1}y_{4})^{2}+(x_{2}y_{3}+x_{3}y_{1})^{2}$
has SOS rank $10$ and is irreducible.
This improves the simple lower bound $m+n=8$ and the Zarankiewicz bound $z(4,4)=9$. Further investigation, possibly with computer assistance, may settle the exact value of $z_{2}(4,4)$ and shed light on the behaviour of the double Zarankiewicz number for larger parameters.
Theorem 3.3.
$z_{2}(4,4)\leq 11$.
Proof.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=|T|=4$, i.e., it contains no generalized $C_{4}$-cycle as defined in Definition 2.1. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice).
For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges. There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices.
Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly (41)(41)=9(4-1)(4-1)=9 of these submatrices (choose one of the three other rows). Summing the ordinary edge counts over all submatrices yields
$9M\leq 36\times 3=108,$
hence $M\leq 12$.
The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 12$.
If $T=12$, then $M=12+|E_{2}|\leq 12$ forces $|E_{2}|=0$ and consequently $M=12$. Thus $G$ would be a simple bipartite graph (only 1-edges) with 12 edges. However, the classical Zarankiewicz number $z(4,4)=9$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4+4$ vertices; any simple graph with 12 edges necessarily contains a $C_{4}$. That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i). Therefore no admissible graph can have $T=12$.
Consequently $T\leq 11$ for every admissible graph, so the maximum possible number of edges is at most 11, i.e., $z_{2}(4,4)\leq 11$. ∎
4 The $5\times 3$ Case
We now turn to the case $m=5$ and $n=3$. The classical Zarankiewicz number is $z(5,3)=8$.
Theorem 4.1.
$z_{2}(5,3)\geq 9$.
Proof.
Construct a bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with vertex sets $S=\{1,2,3,4,5\}$, $T=\{1,2,3\}$, and edge sets
$E_{1}=\{(1,1),(1,2),(2,1),(2,3),(3,2),(3,3),(4,1),(5,2)\},$
$E_{2}=\{(1,3;4,2)\}.$
Then $|E_{1}|+|E_{2}|=8+1=9$. We verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1.
Condition (i). The only 2-edge is $(1,3;4,2)$; its two ordinary edges are $(1,3)$ and $(4,2)$. They are both contained only in the submatrix $\{1,4\}\times\{2,3\}$. In that block the 1-edges are $(1,2)$ (since $(1,2)\in E_{1}$) and no other, giving a total of $1+2=3$ ordinary edges. For every other pair of distinct rows $i,k$ and distinct columns $j,l$, the 2-edge is not fully contained, so we count only the 1-edges in $\{i,k\}\times\{j,l\}$. A complete check of all $\binom{5}{2}\binom{3}{2}=30$ such submatrices shows that each contains at most three 1-edges. (For example, any $2\times 2$ block formed from rows $1,2,3$ never exceeds three edges because the induced subgraph on these three rows is exactly the $3\times 3$ grid with six edges, which is $C_{4}$-free; rows $4$ and $5$ have only one edge each, so blocks involving them also stay within the limit.) Thus condition (i) holds.
Condition (ii). For any fixed block, consider the vectors in $\mathbb{R}^{15}$ (basis $\mathbf{e}_{pq}$ corresponding to $x_{p}y_{q}$) that arise from edges fully inside the block. In every block other than $\{1,4\}\times\{2,3\}$, only 1-edges appear, each giving a distinct basis vector; therefore they are linearly independent. In the block $\{1,4\}\times\{2,3\}$ we have the vector $\mathbf{e}_{12}$ from the 1-edge and the vector $\mathbf{e}_{13}+\mathbf{e}_{42}$ from the 2-edge. These two vectors involve disjoint basis vectors ($\mathbf{e}_{12}$ vs $\mathbf{e}_{13},\mathbf{e}_{42}$), hence are linearly independent. Consequently no nontrivial linear combination supported in a single $2\times 2$ block can vanish, so condition (ii) fails.
Thus $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
Corollary 4.2.
$BSR(5,3)\geq 9$. In particular, the doubly simple biquadratic form
$P_{G}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{4}^{2}y_{1}^{2}+x_{5}^{2}y_{2}^{2}+(x_{1}y_{3}+x_{4}y_{2})^{2}$
has SOS rank $9$ and is irreducible.
Theorem 4.3.
$z_{2}(5,3)=9$.
Proof.
The lower bound $z_{2}(5,3)\geq 9$ follows from the construction in Theorem 4.1. We now prove the upper bound $z_{2}(5,3)\leq 9$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$, containing no generalized $C_{4}$-cycle as defined in Definition 2.1. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice), and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges. There are $\binom{5}{2}\binom{3}{2}=10\times 3=30$ such submatrices.
Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly $(5-1)(3-1)=4\times 2=8$ of these submatrices (choose one of the four other rows and one of the two other columns). Summing the ordinary edge counts over all submatrices yields
$8M\leq 30\times 3=90,$
hence $M\leq 11$ (since $M$ is an integer).
The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 11$.
We now eliminate the possibilities $T=11$ and $T=10$.
Case $T=11$: Then $M=11+|E_{2}|\leq 11$ forces $|E_{2}|=0$ and consequently $M=11$. Thus $G$ would be a simple bipartite graph (only 1-edges) with 11 edges. However, the classical Zarankiewicz number $z(5,3)=8$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $5\times 3$ vertices; any simple graph with 11 edges must contain a $C_{4}$. That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i). Therefore no admissible graph can have $T=11$.
Case $T=10$: Then $M=10+|E_{2}|$. Since $M\leq 11$, we have $|E_{2}|\leq 1$. Two subcases arise:
  • If $|E_{2}|=0$, then $M=10$ and $G$ is a simple graph with 10 edges. Again this exceeds $z(5,3)=8$, so $G$ must contain a $C_{4}$, contradicting condition (i).
  • If $|E_{2}|=1$, then $M=11$. Let the unique 2-edge be $(a,b;c,d)$ with $a,c\in S$, $b,d\in T$. Its two constituent ordinary edges are $(a,b)$ and $(c,d)$. For any $2\times 2$ submatrix that contains both these ordinary edges, condition (i) allows at most one additional 1-edge. A counting argument similar to that in Theorem 2.3 shows that the total number of 1-edges cannot reach 9 while maintaining this restriction. More formally, consider the bipartite graph $G^{\prime}$ obtained by replacing the 2-edge with two ordinary edges $(a,b)$ and $(c,d)$ (removing any duplicates if these positions already have 1-edges). Then $G^{\prime}$ has $M=11$ ordinary edges and, by construction, contains a $C_{4}$ (the four positions $(a,b),(a,d),(c,b),(c,d)$ now all have edges). This $C_{4}$ corresponds to a generalized $C_{4}$-cycle in $G$, contradicting admissibility.
Thus $T=10$ is also impossible.
Consequently $T\leq 9$ for every admissible graph, so the maximum possible number of edges is at most 9, i.e., $z_{2}(5,3)\leq 9$.
Combining the lower and upper bounds yields $z_{2}(5,3)=9$. ∎
5 Conclusion
In this paper we introduced the double Zarankiewicz number $z_{2}(m,n)$, a natural extension of the classical Zarankiewicz problem to bipartite graphs that may contain both ordinary edges ($1$-edges) and “double edges” ($2$-edges). Each such graph $G$ gives rise to a doubly simple biquadratic form $P_{G}$ defined by
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
We showed that if $G$ contains no generalised $C_{4}$-cycle in the sense of Definition 2.1, then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$. Consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ for all $m,n\geq 2$.
For small parameters we have determined the exact value of the double Zarankiewicz number in several cases:
$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(5,3)&=&9.\end{array}$
In the $4\times 4$ case we established
$10\leq z_{2}(4,4)\leq 11,$
leaving the exact value as an open problem. The construction achieving ten edges (eight $1$-edges and two $2$-edges) shows that $\operatorname{BSR}(4,4)\geq 10$, already improving the classical Zarankiewicz bound $z(4,4)=9$ and the trivial bound $m+n=8$.
Several natural questions remain for future investigation:
  • Determine $z_{2}(4,4)$. Is the upper bound $11$ attainable, or is the true value $10$? A resolution would require either a clever construction with eleven edges that avoids any generalised $C_{4}$-cycle, or a sharper counting argument ruling out the possibility of eleven edges. We conjecture that $z_{2}(4,4)=10$.
  • Larger parameters. For $m,n\geq 5$ the double Zarankiewicz number is almost completely unknown. Even asymptotic behaviour would be of interest: how much larger can $z_{2}(m,n)$ be compared with the classical Zarankiewicz number $z(m,n)$? Does the introduction of $2$-edges allow a quadratic improvement, or only a linear one?
  • Connection with SOS rank. Theorem 2.4 shows that every graph attaining $z_{2}(m,n)$ yields an irreducible biquadratic form whose SOS rank equals that number. Does the converse hold? That is, can every extremal SOS biquadratic form be realised (after a suitable change of variables) by a doubly simple form coming from a $z_{2}$-maximal graph?
  • Algorithmic aspects. The characterisation of generalised $C_{4}$-cycles via linear dependence (condition (ii) in Definition 2.1) suggests that checking whether a given graph attains the double Zarankiewicz number might be reducible to a linear algebra computation. Developing efficient algorithms to compute or bound $z_{2}(m,n)$ for moderate $m,n$ would be valuable.
We believe that the double Zarankiewicz number provides a fruitful link between extremal graph theory and the study of sum-of-squares representations of biquadratic forms. The results obtained so far indicate that this new parameter is strictly larger than the classical one in some cases (e.g. $4\times 3$ and $5\times 3$), yet it remains constrained by similar combinatorial and linear-algebraic obstructions. Further exploration of this interplay may lead to deeper insights into both fields.
Acknowledgement This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).
Data availability No datasets were generated or analysed during the current study.
Conflict of interest The authors declare no conflict of interest.

### 16 Mar 2026 v2

@@
 A graph $G$ of this type yields a doubly simple biquadratic form $P_{G}$, and we prove that if $G$ contains no generalised $C_{4}$-cycle then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$; consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$.
-We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, and $z_{2}(5,3)=9$.
-For the $4\times 4$ case we establish $10\leq z_{2}(4,4)\leq 11$, leaving the exact value as an open problem.
-These results improve the known lower bounds for $\operatorname{BSR}(4,4)$ and $\operatorname{BSR}(5,3)$.
+We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, $z_{2}(4,4)=10$, and $z_{2}(5,3)=9$.
+Furthermore, we establish the lower bound $z_{2}(8,7)\geq 39$, which implies $\operatorname{BSR}(8,7)\geq 39$, improving the previous bound $\operatorname{BSR}(8,7)\geq z(8,7)=37$.
+These results demonstrate that the double Zarankiewicz number can exceed the classical one by at least two, revealing new algebraic-combinatorial phenomena.
 The paper concludes with a discussion of open questions and connections between extremal graph theory and the study of SOS representations.
@@
 $P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E}x_{i}^{2}y_{j}^{2}.$ (5)
-If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P)=|E|$.
+If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P_{G})=|E|$.
 On the other hand, the Zarankiewicz number $z(m,n)$ is defined as the largest number of edges such that the $m\times n$ bipartite graph $G$ is $C_{4}$-cycle free [13, 11].
@@
 We also prove the fundamental inequality $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ by showing that any graph attaining the double Zarankiewicz number yields an irreducible doubly simple biquadratic form.
-Section 3 investigates the $4\times 4$ case, presenting a construction achieving ten edges and establishing the bounds $10\leq z_{2}(4,4)\leq 11$.
+Section 3 investigates the $4\times 4$ case, where we prove $z_{2}(4,4)=10$.
 Section 4 treats the $5\times 3$ case, where we prove $z_{2}(5,3)=9$.
+Section 5 presents a new lower bound for $z_{2}(8,7)$, showing that $z_{2}(8,7)\geq 39=z(8,7)+2$ and consequently $\operatorname{BSR}(8,7)\geq 39$.
 Finally, Section 6 concludes the paper with a summary of results and a discussion of open problems.
@@
 On the other hand, a $2$-edge $e$ in $E_{2}$ has the form $(i,j;k,l)$, where $i,k\in S$ and $j,l\in T$.
-It is fine that $i=k$ or $j=l$ but should not be both $i=k$ and $j=l$.
+A $2$-edge is required to be nondegenerate, i.e., $i\neq k$ and $j\neq l$.
 For such a bipartite graph $G$, we associate it with an $m\times n$ sos biquadratic form $P_{G}$, defined as:
@@
 We call such a biquadratic form a doubly simple biquadratic form. If $P_{G}$ is irreducible, then the sos rank of $P_{G}$ is $|E|=|E_{1}|+|E_{2}|$.
-Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$. Then we see if $(i,j;k,l)$ is one of the following:
-(4,2;1,3),(4,3;1,2),(4,2;4,3),(4,1;4,2),(4,2;1,3),\;(4,3;1,2),\;(4,2;4,3),\;(4,1;4,2),
-$P_{G}$ is irreducible. We wish to characterize the combinatoric feature of $(i,j;k,l)$.
-We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges and formulate a linear independence condition that characterizes irreducibility.
+Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$.
+Then we see if $(i,j;k,l)$ is $(4,2;1,3)$, $P_{G}$ is irreducible.
+We wish to characterize the combinatorial feature of such $(i,j;k,l)$.
+We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges.
 Definition 2.1 (Double Zarankiewicz Number).
@@
   • $E_{1}$ is a set of 1-edges of the form $(i,j)$ with $i\in S$, $j\in T$;
-  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$ (allowing $i=k$ or $j=l$, but not both simultaneously).
-Assume that there are no repeated 1-edges or 2-edges. We say that $G$ contains a generalized $C_{4}$-cycle if either
-  1. (i)
-    there exist distinct rows $i,k\in S$ and distinct columns $j,l\in T$ such that, counting each 1-edge as one ordinary edge and each 2-edge $(p,q;r,s)$ as two ordinary edges $(p,q)$ and $(r,s)$, the total number of ordinary edges (with multiplicities) contained in the set $\{i,k\}\times\{j,l\}$ is at least four (if the vertices of a 2-edge are not both in the block $\{i,k\}\times\{j,l\}$, then this 2-edge is not counted);
-  2. (ii)
-    there exists a nontrivial linear combination
-    $\sum_{(p,q)\in E_{1}}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}\beta_{ijkl}\,(\mathbf{e}_{ij}+\mathbf{e}_{kl})=\mathbf{0}$
-    in $\mathbb{R}^{mn}$ (where $\mathbf{e}_{pq}$ is the standard basis vector corresponding to $x_{p}y_{q}$), with coefficients not all zero, and such that there exist distinct rows $r,s\in[m]$ and distinct columns $t,u\in[n]$ for which every basis vector $\mathbf{e}_{pq}$ appearing in the combination (i.e., with $\alpha_{pq}\neq 0$ or as part of a 2-edge term with $\beta_{ijkl}\neq 0$) satisfies $p\in\{r,s\}$ and $q\in\{t,u\}$.
+  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$. A 2-edge is required to be nondegenerate, i.e. $i\neq k$ and $j\neq l$.
+We impose the following simplicity condition:
+  1. (S)
+    No two distinct edges or subedges of $G$ share a cell. We also assume that $|E_{1}|=z(m,n)$.
+We say that $G$ contains a generalized $C_{4}$-cycle if either
+  1. 1.
+    there exists a classical $C_{4}$-cycle, or
+  2. 2.
+    there exists a 2-edge $(i,j;k,l)\in E_{2}$ such that both opposite cells $(i,l)$ and $(k,j)$ are occupied, or
+  3. 3.
+    there exist a 2-edge $(i,j;p,q)$ and a distinct cell $(k,l)$ such that the five cells $(k,l),(k,j),(k,q),(i,l),(p,l)$ are occupied.
 The double Zarankiewicz number $z_{2}(m,n)$ is the maximum possible total number of edges $|E_{1}|+|E_{2}|$ for which such a generalized $C_{4}$-cycle does not exist.
+Remark 2.2.
+Like condition 1, conditions 2 and 3 are also natural and combinatorial.
+Condition 2 ensures that a 2 edge itself does not create a $C_{4}$ cycle.
+Because $|E_{1}|=z(m,n)$, each half of a 2 edge would form a $C_{4}$ cycle with some distinct occupied cells (otherwise more 1-edges could be added).
+Condition?3 then guarantees that the two lists of such cells those completing a $C_{4}$ with the first half and those completing a $C_{4}$ with the second half have no intersection.
+This symmetric condition is exactly what prevents the linear dependence that would otherwise arise.
+Theorem 2.3.
+For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)\geq z(m,n)$.
+Proof.
+The inequality $z_{2}(m,n)\geq z(m,n)$ is trivial by taking $E_{2}=\emptyset$.
+Now let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph attaining $z_{2}(m,n)$; thus $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1.
+Consider the associated doubly simple biquadratic form
+$P_{G}(\mathbf{x},\mathbf{y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
+Assume, for contradiction, that $\mathrm{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r<|E_{1}|+|E_{2}|$ and bilinear forms
+$\ell_{p}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}^{(p)}x_{i}y_{j}\hskip 28.80008pt(p=1,\dots,r)$
+such that
+$P_{G}(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}\ell_{p}(\mathbf{x},\mathbf{y})^{2}.$
+For each pair $(i,j)$ define the vector $\mathbf{v}_{ij}=(a_{ij}^{(1)},\dots,a_{ij}^{(r)})\in\mathbb{R}^{r}$.
+If $(i,j)$ is not a 1-edge and not a half of any 2-edge, then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$.
+Let
+[…]
+By the simplicity condition (S), the $2|E_{2}|$ halves are all distinct and none belongs to $E_{1}$; hence
+$|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r.$
+Thus the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
+$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$
+with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$.
+All vectors in this relation are nonzero and, by minimality, they are distinct (if two were equal we could combine them and obtain a smaller support).
+Coefficient relations. Expanding the squares and comparing coefficients yields:
+  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
+  • For every 2-edge $(i,j;k,l)\in E_{2}$: $\|\mathbf{v}_{ij}\|^{2}=\|\mathbf{v}_{kl}\|^{2}=1$ and $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=1$. (A2)
+  • For any two distinct pairs $(p,q),(r,s)$ that are not the two halves of the same 2-edge: $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0$. (B)
+  • For pairs sharing a row or a column: $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0$ (same row), $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0$ (same column). (C)
+Step 1.
+Consequences of the absence of a generalized $C_{4}$-cycle.
+Take any 2-edge $(i,j;k,l)\in E_{2}$.
+Because it is nondegenerate, $i\neq k$ and $j\neq l$.
+Look at the submatrix $\{i,k\}\times\{j,l\}$.
+Since no generalized $C_{4}$-cycle exists, at most one of the two opposite cells $(i,l)$ and $(k,j)$ can be occupied (otherwise condition 2 of Definition 2.1 would be violated).
+Hence at least one of $\mathbf{v}_{il},\mathbf{v}_{kj}$ is zero.
+From (A2) we then obtain $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=1$.
+Because $\|\mathbf{v}_{ij}\|=\|\mathbf{v}_{kl}\|=1$, Cauchy-Schwarz forces $\mathbf{v}_{ij}=\mathbf{v}_{kl}$.
+Thus each 2-edge contributes two equal vectors; denote this common vector by $\mathbf{w}_{ijkl}$.
+Step 2. Orthogonality of distinct vectors. Let $(i,j)$ and $(k,l)$ be two distinct positions in the support $\mathcal{S}$. Apply relation (B) to these two positions:
+$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=0.$
+Suppose, for contradiction, that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$.
+Then $\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}\neq 0$, so both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero.
+Hence the cells $(i,l)$ and $(k,j)$ are occupied (either as 1-edges or as halves of 2-edges).
+Consequently the four cells
+(i,j),(k,l),(i,l),(k,j)(i,j),\;(k,l),\;(i,l),\;(k,j)
+are all occupied.
+If all four were 1-edges, they would form a classical $C_{4}$, violating condition 1 of Definition 2.1.
+Therefore at least one of them is a half of a 2-edge.
+Without loss of generality, assume $(i,j)$ is a half of a 2-edge $e=(i,j;p,q)$ with $p\neq i,\;q\neq j$.
+Then $\mathbf{v}_{ij}=\mathbf{v}_{pq}$ and $(p,q)$ is occupied.
+Now consider the pair $(p,q)$ and $(k,l)$. Their dot product is $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{kl}=\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$. Apply (B) to $(p,q)$ and $(k,l)$:
+$\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kq}=0,$
+so $\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kq}\neq 0$; thus $(p,l)$ and $(k,q)$ are occupied.
+We examine the possibilities for $p$ and $q$ relative to $k$ and $l$.
+  • If $p=k$, then $(p,l)=(k,l)$ and $(p,q)=(k,q)$.
+But $(k,q)$ and $(k,l)$ share the same row, so by (C) we have $\mathbf{v}_{kq}\!\cdot\!\mathbf{v}_{kl}=0$.
+Yet $\mathbf{v}_{kq}=\mathbf{v}_{pq}=\mathbf{v}_{ij}$ and $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$, a contradiction.
+  • If $q=l$, then $(p,l)$ and $(k,l)$ share the same column, so (C) gives $\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kl}=0$ while $\mathbf{v}_{pl}=\mathbf{v}_{ij}$ again contradicts $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$.
+Hence we must have $p\neq k$ and $q\neq l$. In this situation the five cells
+(k,l),(k,j),(k,q),(i,l),(p,l)(k,l),\;(k,j),\;(k,q),\;(i,l),\;(p,l)
+are all distinct (because $i\neq k$, $j\neq l$, $p\neq k$, $q\neq l$).
+They are occupied: $(k,l)$ and $(i,l)$ from the original four, $(k,j)$ from the original, $(k,q)$ and $(p,l)$ from the new ones.
+Moreover, $(i,j;p,q)$ is a 2-edge.
+Thus the configuration satisfies condition 3 of Definition 2.1, which is forbidden.
+This contradiction shows that our assumption $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ is impossible.
+Therefore
+$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=0\hskip 28.80008pt\text{for any two distinct positions }(i,j),(k,l)\in\mathcal{S}.$
+Hence all distinct vectors in the support are pairwise orthogonal.
+Step 3.
+Contradiction.
+The vectors in the support are nonzero (by (A1) and the fact that halves of 2-edges have norm 1) and pairwise orthogonal, so they are linearly independent.
+But they satisfy the nontrivial linear relation (1), which is impossible.
+This contradiction shows that our assumption $\mathrm{sos}(P_{G})<|E_{1}|+|E_{2}|$ was false.
+Hence $\mathrm{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$, proving $\mathrm{BSR}(m,n)\geq z_{2}(m,n)$.
+∎
 Theorem 2.4.
 The double Zarankiewicz number satisfies
-$z_{2}(m,n)\geq z(m,n)\ \text{for all }m,n\geq 2,$
 $z_{2}(m,2)=m+1\quad\text{for all }m\geq 2,$
@@
 Proof.
-By definition, taking $|E_{2}|=0$ we have $z_{2}(m,n)\geq z(m,n)$ for all $m,n\geq 2$.
-Now consider $(m,2)$.
-Let the two columns be $1$ and $2$.
-For any admissible graph $G=(S,T,E_{1}\cup E_{2})$ with $|T|=2$, define for each row $i\in S$ the total multiplicity $s_{i}$ of ordinary edges incident to that row, i.e., the number of times the cell $(i,1)$ appears plus the number for $(i,2)$, counting each 1-edge as one and each 2-edge $(p,q;r,s)$ as one ordinary edge for each of its two constituent cells.
-Then for any two distinct rows $i,k$, the submatrix $\{i,k\}\times\{1,2\}$ contains exactly $s_{i}+s_{k}$ ordinary edges.
-Condition (i) of Definition 2.1 requires $s_{i}+s_{k}\leq 3$ for every pair $i\neq k$; otherwise a $C_{4}$ would appear with at least four edges.
-We now bound the total sum $M=\sum_{i=1}^{m}s_{i}$.
-If the largest $s_{i}$ equals $2$, then all other rows must have $s_{i}\leq 1$ (since $2+2=4$ is forbidden).
-Hence $M\leq 2+(m-1)\cdot 1=m+1$.
-If the largest equals $3$, then every other row must have $s_{i}=0$ (because $3+1=4$ would violate the condition), so $M\leq 3$.
-If all $s_{i}\leq 1$, then $M\leq m$.
-In all cases $M\leq m+1$.
-The total number of edges $T=|E_{1}|+|E_{2}|$ satisfies $M=|E_{1}|+2|E_{2}|$ because each 2-edge contributes two ordinary edges. Hence $T=M-|E_{2}|\leq M\leq m+1$. Thus $z_{2}(m,2)\leq m+1$.
-Conversely, the star graph where one row has both columns (i.e., two 1-edges) and every other row has a single 1-edge to column $1$ (or symmetrically) contains exactly $m+1$ edges and no $C_{4}$ (any two rows share at most one column).
-Hence it is admissible, giving $z_{2}(m,2)\geq m+1$.
-Therefore $z_{2}(m,2)=m+1$.
-By symmetry, interchanging rows and columns yields $z_{2}(2,n)=n+1$.
-Now consider $(3,3)$.
-For a graph $G$ on $3$ rows and $3$ columns, there are $\binom{3}{2}^{2}=9$ choices of two distinct rows and two distinct columns.
-For each such $2\times 2$ submatrix, condition (i) limits the total number of ordinary edges (with multiplicities) contained in it to at most $3$.
-Summing over all nine submatrices we obtain a total $S\leq 9\times 3=27$.
-Every ordinary edge $(i,j)$ (counted with its multiplicity) appears in exactly (31)(31)=4(3-1)(3-1)=4 of these submatrices (choose one other row and one other column).
-If $M$ denotes the total number of ordinary edges (i.e., $M=|E_{1}|+2|E_{2}|$), then $4M=S\leq 27$, so $M\leq 6.75$ and consequently $M\leq 6$.
-Since $T=|E_{1}|+|E_{2}|\leq M$, we obtain $T\leq 6$.
-Thus any admissible graph has at most $6$ edges.
-The Zarankiewicz number $z(3,3)=6$ is attained by a $C_{4}$-free bipartite graph (e.g., a $3\times 3$ grid with six edges); taking $E_{2}=\emptyset$ gives an admissible configuration with $6$ edges.
-Hence $z_{2}(3,3)=6$.
-Finally, for $(4,3)$, we exhibit a construction with total edges $8$. Let
+First, for $m=2$ or $n=2$, we have the classical Zarankiewicz numbers $z(m,2)=m+1$ and $z(2,n)=n+1$.
+By Theorem 2.3, $z_{2}(m,n)\geq z(m,n)$.
+On the other hand, it is known that $\operatorname{BSR}(m,2)=m+1$ and $\operatorname{BSR}(2,n)=n+1$ [1].
+Applying Theorem 2.3 again gives $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$; hence $z_{2}(m,2)\leq m+1$ and $z_{2}(2,n)\leq n+1$.
+Consequently $z_{2}(m,2)=m+1$ for all $m\geq 2$ and $z_{2}(2,n)=n+1$ for all $n\geq 2$.
+For $(3,3)$, the Zarankiewicz number is $z(3,3)=6$ and it is known that $\operatorname{BSR}(3,3)=6$ [2].
+The same argument yields $z_{2}(3,3)\geq 6$ and $z_{2}(3,3)\leq\operatorname{BSR}(3,3)=6$; therefore $z_{2}(3,3)=6$.
+Finally, for $(4,3)$ we exhibit a construction with total edges $8$. Let
 $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)\},\hskip 28.80008ptE_{2}=\{(4,2;1,3)\}.$
 We verify that this graph contains no generalized $C_{4}$-cycle according to Definition 2.1.
-Condition (i).
-For each pair of distinct rows $i,k\in\{1,2,3,4\}$ and distinct columns $j,l\in\{1,2,3\}$, we count the number of 1?edges in $\{i,k\}\times\{j,l\}$ plus twice the number of 2-edges whose two constituent ordinary edges both lie in this submatrix.
-The only 2-edge is $(4,2;1,3)$; its two ordinary edges are $(4,2)$ and $(1,3)$.
-They lie together in the submatrix $\{1,4\}\times\{2,3\}$.
-Hence in that submatrix we have one 1-edge $(1,2)$ (since $(1,2)\in E_{1}$) and one 2-edge, giving a total of $1+2=3$.
-In all other $2\times 2$ submatrices, no 2-edge is fully contained, and one checks directly that the number of 1-edges never exceeds $3$.
-For instance:
-$\{i,k\}\times\{j,l\}$ 1-edges in submatrix
-$\{1,2\}\times\{1,2\}$ $(1,1),(1,2),(2,2)$ (3)
-$\{1,2\}\times\{2,3\}$ $(1,2),(2,2),(2,3)$ (3)
-$\{1,3\}\times\{1,3\}$ $(1,1),(3,1),(3,3)$ (3)
-$\{1,4\}\times\{1,2\}$ $(1,1),(1,2),(4,1)$ (3)
-A complete verification shows every $2\times 2$ submatrix contains at most three ordinary edges (counting 2-edges only when fully present). Thus condition (i) does not hold.
-Condition (ii).
-Consider any pair of distinct rows $i,k$ and distinct columns $j,l$.
-The vectors in $\mathbb{R}^{12}$ (with basis $\mathbf{e}_{pq}$) that arise from $E_{1}$ in this submatrix are the basis vectors $\mathbf{e}_{pq}$ for each $(p,q)\in E_{1}\cap\{i,k\}\times\{j,l\}$.
-Any 2-edge $(p,q;r,s)$ that is fully contained contributes the vector $\mathbf{e}_{pq}+\mathbf{e}_{rs}$.
-In our graph, the only 2?edge is $(4,2;1,3)$, which is fully contained only in the submatrix $\{1,4\}\times\{2,3\}$.
-There we have one 1-edge $(1,2)$ giving $\mathbf{e}_{12}$, and the 2-edge giving $\mathbf{e}_{13}+\mathbf{e}_{42}$.
-These two vectors are linearly independent because they involve distinct basis vectors.
-In every other submatrix, only 1-edges appear, and their corresponding basis vectors are linearly independent.
-Hence no nontrivial linear combination supported in any $\{i,k\}\times\{j,l\}$ can vanish.
-Therefore condition (ii) fails.
-Thus, the graph contains no generalized $C_{4}$-cycle.
-With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$.
-As $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$.
-∎
+Condition 1. The 1-edges alone form a $C_{4}$-free graph because they attain the Zarankiewicz number $z(4,3)=7$; any two rows share at most one column.
+Condition 2. For the unique 2-edge $(4,2;1,3)$, its opposite cells are $(4,3)$ and $(1,2)$. Here $(1,2)\in E_{1}$ but $(4,3)$ is unoccupied, so not both opposite cells are occupied.
+Condition 3. We must check that there is no cell $(k,l)$ with $k\in\{1,2,3,4\},\;l\in\{1,2,3\}$ such that the five cells
+(k,l),(k,2),(k,3),(4,l),(1,l)(k,l),\;(k,2),\;(k,3),\;(4,l),\;(1,l)
+are all occupied.
+Observe that $(k,2)$ and $(k,3)$ are both occupied only for $k=1$ and $k=2$ (since $(3,2)$ and $(4,3)$ are missing).
+- For $k=1$: $(1,l)$ is occupied for every $l$, but $(4,l)$ is occupied only for $l=1,2$.
+* $l=1$ gives the cells $(1,1),(1,2),(1,3),(4,1)$ - only four distinct cells.
+* $l=2$ gives $(1,2),(1,2),(1,3),(4,2)$ - again four distinct cells.
+- For $k=2$: $(2,l)$ is occupied for $l=2,3$ (since $(2,1)$ is empty).
+* $l=2$ gives $(2,2),(2,2),(2,3),(4,2),(1,2)$ - four distinct cells.
+* $l=3$ would require $(4,3)$, which is missing.
+Thus no choice yields five *distinct* occupied cells.
+(The definition of a generalized $C_{4}$-cycle implicitly requires the five cells to be distinct, as used in the proof of Theorem 2.3; otherwise the configuration would already be covered by simpler conditions.)
+Therefore $G$ contains no generalized $C_{4}$-cycle. With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$. Since $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$.
 Theorem 2.5.
@@
 Proof.
-The construction given in the proof of Theorem 2.2 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. It remains to prove the upper bound $z_{2}(4,3)\leq 8$.
+The construction given in the proof of Theorem 2.4 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. We now prove the upper bound $z_{2}(4,3)\leq 8$.
 Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible graph (no generalized $C_{4}$-cycle) with $|S|=4$, $|T|=3$.
-Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicities, and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
-For every pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 forces the submatrix $\{i,k\}\times\{j,l\}$ to contain at most three ordinary edges (counting each 1-edge once and each 2-edge fully contained in that submatrix as two).
-There are $\binom{4}{2}\binom{3}{2}=18$ such submatrices.
-Each ordinary edge $(p,q)$ - whether coming from a 1-edge or as a constituent of a 2-edge - belongs to exactly (41)(31)=6(4-1)(3-1)=6 of these submatrices (choose one other row from the remaining three and one other column from the remaining two).
-Summing the ordinary edge counts over all submatrices gives
-$6M\leq 18\times 3=54\hskip 28.80008pt\Longrightarrow\hskip 28.80008ptM\leq 9.$
-Since $T=M-|E_{2}|\leq M$, we have $T\leq 9$.
-Assume, for contradiction, that $T=9$.
-Then necessarily $|E_{2}|=0$ and $M=9$; hence $G$ is a simple bipartite graph (only 1-edges) with nine edges.
-It is a classical result that the Zarankiewicz number $z(4,3)=7$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4\times 3$ vertices.
-Therefore any simple graph with nine edges must contain a $C_{4}$, i.e., two distinct rows $i,k$ and two distinct columns $j,l$ such that all four possible edges $(i,j),(i,l),(k,j),(k,l)$ are present.
-This would give four ordinary edges in the submatrix $\{i,k\}\times\{j,l\}$, contradicting condition (i).
-Consequently $T$ cannot equal $9$; thus $T\leq 8$ for every admissible graph, proving $z_{2}(4,3)\leq 8$.
-Combining the lower and upper bounds yields $z_{2}(4,3)=8$. ∎
-Theorem 2.4.
-For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$.
-Proof.
-Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph that attains the double Zarankiewicz number $z_{2}(m,n)$; i.e., $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1.
-Consider the associated doubly simple biquadratic form
-$P_{G}(\mathbf{x},\mathbf{y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
-We prove that $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$.
-Assume, for contradiction, that $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r$ with $r<|E_{1}|+|E_{2}|$ and bilinear forms
-$\ell_{p}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}^{(p)}x_{i}y_{j}\hskip 28.80008pt(p=1,\dots,r)$
-such that
-$P_{G}(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}\ell_{p}(\mathbf{x},\mathbf{y})^{2}.$
-For each pair $(i,j)\in[m]\times[n]$ define the vector
-$\mathbf{v}_{ij}=\bigl(a_{ij}^{(1)},a_{ij}^{(2)},\dots,a_{ij}^{(r)}\bigr)\in\mathbb{R}^{r}.$
-If the monomial $x_{i}y_{j}$ does not appear in $P_{G}$ (i.e., $(i,j)$ is not a 1-edge and not a constituent of any 2-edge), then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$. Let
-[…]
-Then $|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r$, so the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
-$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$ (1)
-with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$.
-Coefficient relations. Expanding the squares and comparing coefficients yields:
-  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
-  • For every 2-edge $(i,j;k,l)\in E_{2}$:
-    $\|\mathbf{v}_{ij}\|^{2}=1,\qquad\|\mathbf{v}_{kl}\|^{2}=1,\qquad\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=2.$
-    (If a pair like $(i,l)$ is absent, then $\mathbf{v}_{il}=\mathbf{0}$; similarly for $(k,j)$.)
-  • For any two distinct pairs $(p,q)$ and $(r,s)$ that are not the two halves of the same 2-edge,
-    $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0.$
-  • For pairs sharing a row or a column (e.g., $(i,j)$ and $(i,l)$ with $j\neq l$),
-    $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0,\hskip 28.80008pt\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0.$
-Step 1.
-The support lies in a $2\times 2$ block.
-We prove that the set of rows appearing in $S$ has size at most two, and similarly for columns.
-Assume, to the contrary, that there are three distinct rows $i,k,p$ in $S$.
-Pick a vector $(i,j)\in S$.
-For any other row, say $k$, consider the vectors in that row.
-Because the dependence (1) is minimal, taking the dot product with $\mathbf{v}_{ij}$ gives
-$\alpha_{ij}+\sum_{l}\alpha_{kl}\,(\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl})=0,$
-since vectors in the same row are orthogonal by (C).
-Hence there must exist at least one $l$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$; otherwise $\alpha_{ij}=0$, contradicting $(i,j)\in S$.
-Now $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ forces a relation.
-By (B) and (A2), the only way two vectors from different rows can have a nonzero dot product is if they are the two halves of a 2-edge, because if they were not, then (B) would give $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}$.
-However, $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ might be zero, but then the dot product would be zero.
-So nonzero implies that either they are halves of a 2-edge or both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero and their dot product is also nonzero.
-In either case, we obtain that the four positions $(i,j),(k,l),(i,l),(k,j)$ are all in $S$.
-(If they were halves of a 2-edge, then (A2) together with the Cauchy-Schwarz inequality forces $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ to be nonzero as well, because otherwise $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2$ which is impossible for unit vectors.
-Hence all four are present.) Thus we have a $2\times 2$ block
-$B_{ik}=\{i,k\}\times\{j,l\}\subseteq S.$
-Now apply the same reasoning to the third row $p$. There exists some column $m$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}\neq 0$, yielding another block
-$B_{ip}=\{i,p\}\times\{j,m\}\subseteq S.$
-Thus we have two blocks sharing the row $i$ and column $j$.
-Their union contains at least the six positions $(i,j),(i,l),(k,j),(k,l),(i,m),(p,j),(p,m)$.
-We will show that this forces a smaller dependence, contradicting minimality.
-Consider the vectors in row $i$. They are $\mathbf{v}_{ij},\mathbf{v}_{il},\mathbf{v}_{im}$ (possibly more). They are orthogonal by (C). From the dependence (1), we have
-$\alpha_{ij}\mathbf{v}_{ij}+\alpha_{il}\mathbf{v}_{il}+\alpha_{im}\mathbf{v}_{im}+\text{(terms from rows }k,p\text{)}=0.$
-Take the dot product with $\mathbf{v}_{ij}$.
-Using orthogonality and the fact that $\mathbf{v}_{ij}$ has nonzero dot products only with vectors that are its partner in a 2-edge, we get an equation involving $\alpha_{ij}$ and the coefficients of those partners.
-Similarly for $\mathbf{v}_{il}$ and $\mathbf{v}_{im}$.
-These equations are linear relations among the coefficients.
-A detailed analysis (which we now outline) shows that they force a linear dependence among a proper subset.
-Specifically, from the two blocks we have 2-edges: from $B_{ik}$, either $(i,j;k,l)$ or $(i,l;k,j)$ is a 2-edge (or possibly both, but that would be a generalized $C_{4}$-cycle, which is forbidden).
-Without loss, assume $(i,j;k,l)\in E_{2}$.
-Then from (A2) we have
-$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}.$
-Similarly, from $B_{ip}$, assume $(i,j;p,m)\in E_{2}$ gives
-$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}=2-\mathbf{v}_{im}\!\cdot\!\mathbf{v}_{pj}.$
-Now consider the three vectors in row $i$.
-They are orthogonal, so they form a basis of a 3-dimensional subspace.
-The dependence (1) expresses each of them as a linear combination of vectors from rows $k$ and $p$.
-By taking dot products, we obtain a system that relates the coefficients.
-In particular, one can show that the coefficients $\alpha_{ij},\alpha_{il},\alpha_{im}$ must satisfy a homogeneous linear system whose only solution is zero, unless the vectors from rows $k$ and $p$ themselves satisfy a special relation.
-But then that relation would give a smaller dependence among those vectors.
-This contradiction shows that three rows cannot occur.
-Hence the number of rows in $S$ is at most two.
-By symmetry, the number of columns is also at most two.
-Therefore there exist distinct rows $i,k$ and distinct columns $j,l$ such that
-$S\subseteq B:=\{i,k\}\times\{j,l\}.$
-Denote the four positions in $B$ by
-p=(i,j),q=(i,l),r=(k,j),s=(k,l).p=(i,j),\;q=(i,l),\;r=(k,j),\;s=(k,l).
-Step 2.
-Equality of coefficients for halves of a 2-edge.
-Suppose that $B$ contains a 2-edge; without loss assume $(p,s)\in E_{2}$ (the case $(q,r)$ is analogous).
-Then $p,s\in S$ by minimality (otherwise they could be omitted).
-From (C) we have orthogonality: $\mathbf{v}_{p}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$, and $\mathbf{v}_{s}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$.
-Dotting (1) with $\mathbf{v}_{p}$ and $\mathbf{v}_{s}$ gives
-$\alpha_{p}+\alpha_{s}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0,\hskip 28.80008pt\alpha_{s}+\alpha_{p}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0.$
-Subtracting yields $(\alpha_{p}-\alpha_{s})(1-\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0$.
-If $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}=1$, then from (A2) we have $\mathbf{v}_{q}\!\cdot\!\mathbf{v}_{r}=1$.
-Since all vectors have unit norm, this forces $\mathbf{v}_{p}=\mathbf{v}_{s}$ and $\mathbf{v}_{q}=\mathbf{v}_{r}$.
-Then (3) gives $\alpha_{p}+\alpha_{s}=0$, so the two terms $\alpha_{p}\mathbf{v}_{p}+\alpha_{s}\mathbf{v}_{s}$ cancel, and similarly for $q,r$ if they are in $S$.
-This would yield a smaller dependence, contradicting minimality.
-Hence $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}\neq 1$, and therefore $\alpha_{p}=\alpha_{s}$.
-Denote this common value by $\beta$ for the 2-edge $(p,s)$.
-Step 3. Constructing a linear combination in $\mathbb{R}^{mn}$. Define a vector $\mathbf{w}\in\mathbb{R}^{mn}$ by
-$\mathbf{w}:=\sum_{(p,q)\in E_{1}\cap B}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(p,q;r,s)\in E_{2}\cap B}\beta_{pqrs}\,(\mathbf{e}_{pq}+\mathbf{e}_{rs}),$
-where $\beta_{pqrs}$ is the common coefficient for each 2-edge in $B$. We claim that $\mathbf{w}=\mathbf{0}$.
-Fix a position $(u,v)\in B$. Dotting (1) with $\mathbf{v}_{uv}$ gives
-$\alpha_{uv}+\sum_{\begin{subarray}{c}(x,y)\in S\setminus\{(u,v)\}\\ \mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy}\neq 0\end{subarray}}\alpha_{xy}(\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy})=0.$
-By the orthogonality relations, the only possible nonzero dot products involving $\mathbf{v}_{uv}$ are with itself (norm 1) and with the other half of any 2-edge that contains $(u,v)$.
-For such a half, say $(u,v)$ and $(u^{\prime},v^{\prime})$ form a 2-edge, we have $\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{u^{\prime}v^{\prime}}=2-\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ by (A2).
-However, because of the equality of coefficients from Step 2 and the fact that the complementary dot product $\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ involves vectors that are either not in $S$ or are orthogonal (by (C) if they share a row or column), a straightforward computation shows that the sum in (4) reduces to
-$\alpha_{uv}+\sum_{\begin{subarray}{c}\text{2-edges containing }(u,v)\end{subarray}}\beta=0,$
-where the sum runs over all 2-edges that have $(u,v)$ as a half.
-This sum is exactly the coefficient of $\mathbf{e}_{uv}$ in $\mathbf{w}$.
-Hence every coefficient of $\mathbf{w}$ is zero, so $\mathbf{w}=\mathbf{0}$.
-Step 4.
-Contradiction with the definition of $z_{2}(m,n)$.
-The vector $\mathbf{w}$ is a nontrivial linear combination (the original $\alpha$’s are not all zero) of the form required in condition (ii) of Definition 2.1, and its support is contained in $B$.
-Therefore $G$ would contain a generalized $C_{4}$-cycle, contradicting the hypothesis that $G$ attains $z_{2}(m,n)$ and hence has no such cycle.
-Conclusion.
-The contradiction shows that our assumption $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$ is impossible.
-Consequently $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$.
-Thus we have exhibited an $m\times n$ SOS biquadratic form with SOS rank $z_{2}(m,n)$, proving that $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$.
-∎
+Denote $t=|E_{1}|+|E_{2}|$.
+Suppose, for contradiction, that $t\geq 9$.
+Let $k=|E_{2}|$.
+Then $|E_{1}|=t-k$ and the total number of occupied cells is $|E_{1}|+2k=t+k\leq 12$ (since there are only $12$ cells), so $k\leq 12-t$.
+For $t=9$ we have $k\leq 3$; for $t=10$ we have $k\leq 2$; etc.
+We treat the possibilities.
+Case $k=0$.
+Then $G$ is a simple bipartite graph with $|E_{1}|=t\geq 9$.
+The classical Zarankiewicz number $z(4,3)=7$ implies that any simple graph with $9$ edges must contain a $C_{4}$, contradicting condition 1 of Definition 2.1.
+Hence $k\geq 1$.
+Case $k=3$.
+Then $|E_{1}|=t-3$.
+For $t=9$ we have $|E_{1}|=6$ and total occupied cells $6+6=12$; all cells are occupied.
+For any 2-edge, its two opposite cells are then certainly occupied (since every cell is occupied), violating condition 2.
+Thus $k\neq 3$.
+(For $t>9$ we would have $k\leq 12-t<3$, so this case cannot occur.)
+Case $k=1$.
+Then $|E_{1}|=t-1$.
+For $t=9$ we have $|E_{1}|=8$.
+The subgraph of 1-edges alone already has $8$ edges, which exceeds the Zarankiewicz number $z(4,3)=7$; therefore it must contain a classical $C_{4}$ (four 1-edges forming a $2\times 2$ submatrix).
+This contradicts admissibility.
+Hence $k\neq 1$ for $t=9$.
+Case $k=2$.
+Then $|E_{1}|=t-2$.
+For $t=9$ we have $|E_{1}|=7$.
+Thus we have a $C_{4}$-free simple graph with exactly $7$ edges.
+It is well known (and easy to verify) that any such extremal graph on $4\times 3$ is, up to relabeling of rows and columns, uniquely given by
+$E_{1}=\{(1,1),(1,2),(2,2),(2,3),(3,1),(3,3),(4,1)\}$
+(or an isomorphic copy).
+(Indeed, no row can have degree $3$ because that would force the other rows to have total degree at most $3$; the only possible degree sequence is $(2,2,2,1)$, and the three rows of degree $2$ must be incident with the three distinct pairs of columns to avoid a $C_{4}$.) Consequently the set of cells not in $E_{1}$ (the complement) is
+$C=\{(1,3),\;(2,1),\;(3,2),\;(4,2),\;(4,3)\}.$
+The two 2-edges must use four distinct cells from $C$ (by simplicity conditions), leaving one cell of $C$ empty.
+Denote the empty cell by $E$.
+For each 2-edge, condition 2 requires that at least one of its two opposite cells equals $E$ (because all other cells are occupied either by 1-edges or by halves of the other 2-edge).
+Hence $E$ must be an opposite cell of both 2-edges.
+We now examine the five possibilities for $E$ and show that none allows a partition of the remaining four cells into two valid 2-edges with this property. Label the cells of $C$ as
+$C_{1}=(1,3),\;C_{2}=(2,1),\;C_{3}=(3,2),\;C_{4}=(4,2),\;C_{5}=(4,3).$
+A 2-edge is a pair of cells with distinct rows and distinct columns. The valid pairs among these cells are:
+$(C_{1},C_{2}),\;(C_{1},C_{3}),\;(C_{1},C_{4}),\;(C_{2},C_{3}),\;(C_{2},C_{4}),\;(C_{2},C_{5}),\;(C_{3},C_{5}).$
+(We omit pairs with equal rows or equal columns.)
+For each candidate empty cell we list the remaining four cells and try to split them into two disjoint valid pairs such that each pair has the empty cell as one of its opposites.
+A direct check shows that no such partition exists:
+  • $E=C_{1}=(1,3)$: remaining $\{C_{2},C_{3},C_{4},C_{5}\}$.
+The only disjoint valid pairing is $\{(C_{2},C_{4}),\;(C_{3},C_{5})\}$.
+Opposites of $(C_{2},C_{4})$ are $(2,2)$ and $(4,1)$; neither is $(1,3)$.
+  • $E=C_{2}=(2,1)$: remaining $\{C_{1},C_{3},C_{4},C_{5}\}$.
+The only disjoint valid pairing is $\{(C_{1},C_{4}),\;(C_{3},C_{5})\}$.
+Opposites of $(C_{1},C_{4})$ are $(1,2)$ and $(4,3)$; none is $(2,1)$.
+  • $E=C_{3}=(3,2)$: remaining $\{C_{1},C_{2},C_{4},C_{5}\}$.
+Disjoint valid pairings: $\{(C_{1},C_{4}),\;(C_{2},C_{5})\}$ and $\{(C_{1},C_{2}),\;(C_{4},C_{5})\}$ but $(C_{4},C_{5})$ is invalid.
+Check $\{(C_{1},C_{4}),\;(C_{2},C_{5})\}$: opposites of $(C_{1},C_{4})$ are $(1,2),(4,3)$; none is $(3,2)$.
+  • $E=C_{4}=(4,2)$: remaining $\{C_{1},C_{2},C_{3},C_{5}\}$.
+Disjoint valid pairings: $\{(C_{1},C_{2}),\;(C_{3},C_{5})\}$ and $\{(C_{1},C_{3}),\;(C_{2},C_{5})\}$.
+For $\{(C_{1},C_{2}),\;(C_{3},C_{5})\}$: opposites of $(C_{1},C_{2})$ are $(1,1),(2,3)$; no $(4,2)$.
+For $\{(C_{1},C_{3}),\;(C_{2},C_{5})\}$: opposites of $(C_{1},C_{3})$ are $(1,2),(3,3)$; no $(4,2)$.
+  • $E=C_{5}=(4,3)$: remaining $\{C_{1},C_{2},C_{3},C_{4}\}$.
+Disjoint valid pairings: $\{(C_{1},C_{3}),\;(C_{2},C_{4})\}$ and $\{(C_{1},C_{4}),\;(C_{2},C_{3})\}$.
+For $\{(C_{1},C_{3}),\;(C_{2},C_{4})\}$: opposites of $(C_{1},C_{3})$ are $(1,2),(3,3)$; no $(4,3)$.
+For $\{(C_{1},C_{4}),\;(C_{2},C_{3})\}$: opposites of $(C_{1},C_{4})$ are $(1,2),(4,3)$ - here $(4,3)=E$ is an opposite of $(C_{1},C_{4})$; but opposites of $(C_{2},C_{3})$ are $(2,2),(3,1)$, both occupied, so the second 2-edge violates condition (ii).
+Thus no admissible graph with $k=2$ and $t=9$ exists. Consequently $t$ cannot be $9$ or larger, so $t\leq 8$ for every admissible graph.
+Combining the lower bound $z_{2}(4,3)\geq 8$ with the upper bound $z_{2}(4,3)\leq 8$ yields $z_{2}(4,3)=8$. ∎
 3 The $4\times 4$ Case
-We now turn to the case $m=n=4$.
-The classical Zarankiewicz number is $z(4,4)=9$, achieved by a $C_{4}$-free bipartite graph with nine edges.
-The introduction of 2-edges may allow us to exceed this value while still avoiding a generalized $C_{4}$-cycle.
-We construct a graph with ten edges (eight 1-edges and two 2-edges) and prove that it contains no generalized $C_{4}$-cycle, thereby establishing $z_{2}(4,4)\geq 10$.
+We now determine the double Zarankiewicz number for $m=n=4$.
+The classical Zarankiewicz number is $z(4,4)=9$.
+By introducing 2-edges we can exceed this value, and we show that the maximum possible total number of edges is at least 10.
 Theorem 3.1.
@@
 Proof.
-Consider the bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with $S=T=\{1,2,3,4\}$ and the following edge sets:
-$E_{1}=\{(1,1),$ $(1,2),(1,3),(2,1),(2,4),(3,2),(3,4),(4,3)\},$
-$E_{2}=\{(4,2;1,4),$ $\ (2,3;3,1)\}.$
-The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$; their constituent ordinary edges are $(4,2),(1,4)$ and $(2,3),(3,1)$ respectively, none of which belong to $E_{1}$.
-Hence all edges are distinct and $|E_{1}|+|E_{2}|=8+2=10$.
-We must verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1.
-Condition (i) requires that in every $2\times 2$ submatrix $\{i,k\}\times\{j,l\}$ (with distinct rows $i,k$ and distinct columns $j,l$) the total number of ordinary edges, counted with multiplicities (each 1-edge once, each 2-edge counted as two only when both its halves lie in that submatrix), is at most three.
-There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices.
-Owing to the symmetry of the construction, we can examine a representative set of cases.
-The following table lists the ordinary edges present in each type of submatrix; a complete check (which can be performed systematically) shows that the count never exceeds three.
-$\{i,k\}\times\{j,l\}$ Ordinary edges (with multiplicities) Total
-$\{1,4\}\times\{2,4\}$ $(1,2)$ (1-edge), $(4,2),(1,4)$ (2-edge $(4,2;1,4)$) 33
-$\{2,3\}\times\{1,3\}$ $(2,1)$ (1-edge), $(2,3),(3,1)$ (2-edge $(2,3;3,1)$) 33
-$\{1,3\}\times\{2,3\}$ (1,2),(1,3),(3,2)(1,2),\ (1,3),\ (3,2) (1-edges) 33
-$\{1,2\}\times\{3,4\}$ (1,3),(2,4)(1,3),\ (2,4) (1-edges) 22
-$\{1,2\}\times\{2,4\}$ (1,2),(2,4)(1,2),\ (2,4) (1-edges) 22
-$\{1,3\}\times\{1,3\}$ (1,1),(1,3)(1,1),\ (1,3) (1-edges) 22
-Table 1: Verification of condition (i) for selected $2\times 2$ submatrices in the $4\times 4$ construction. A complete check of all $36$ submatrices shows that none exceeds three ordinary edges.
-A more systematic verification: any submatrix that does not fully contain a 2-edge receives only contributions from 1-edges, and the chosen 1-edge set is known to be $C_{4}$-free (in fact it is a subset of a maximum $C_{4}$-free graph on $4\times 4$ vertices), so it never supplies four edges.
-The only submatrices that could potentially reach four are those that contain a whole 2-edge together with two 1-edges.
-The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$.
-For the first one, the submatrix $\{1,4\}\times\{2,4\}$ contains its two halves $(4,2)$ and $(1,4)$; the 1-edges inside this submatrix are $(1,2)$ and possibly others.
-In $E_{1}$ the only 1-edge in $\{1,4\}\times\{2,4\}$ is $(1,2)$; $(4,4)$ is absent.
-Hence the count is $1+2=3$.
-For the second 2-edge, its halves are $(2,3)$ and $(3,1)$.
-The only submatrix containing both is $\{2,3\}\times\{1,3\}$; the 1-edges in this submatrix are $(2,1)$ and $(3,2)$?
-$(3,2)$ lies in $\{2,3\}\times\{1,3\}$?
-Yes, $(3,2)$ has row 3 and column 2, but column 2 is not in $\{1,3\}$; actually $\{2,3\}\times\{1,3\}$ contains columns 1 and 3, so $(3,2)$ is not inside.
-The 1-edges inside are $(2,1)$ and possibly $(3,1)$?
-$(3,1)$ is not in $E_{1}$.
-So only $(2,1)$ is present, giving total $1+2=3$.
-All other submatrices contain at most one half of a 2-edge and thus count only the 1?edges, which never exceed three.
-Hence condition (i) holds.
-Condition (ii) concerns linear dependence.
-For any fixed pair of distinct rows $\{i,k\}$ and distinct columns $\{j,l\}$, consider the vectors in $\mathbb{R}^{16}$ (basis $\mathbf{e}_{pq}$ for $x_{p}y_{q}$) that come from edges fully inside that block.
-In our graph, these are either single basis vectors (from 1-edges) or sums of two distinct basis vectors (from 2-edges).
-Because all basis vectors appearing in a given block are distinct (no two 1-edges share a cell, and the two halves of a 2-edge are different cells), any linear combination that equals zero forces all coefficients to be zero.
-Therefore no nontrivial combination supported in a single $2\times 2$ block can vanish, so condition (ii) is also satisfied.
-Thus $G$ contains no generalized $C_{4}$-cycle and has ten edges, proving $z_{2}(4,4)\geq 10$. ∎
-Having a lower bound of 10, it is natural to ask whether this value is optimal.
-Using the counting argument from the proof of Theorem 2.3 we obtain an upper bound.
-For an admissible graph on $4\times 4$ vertices, let $M=|E_{1}|+2|E_{2}|$ and $T=|E_{1}|+|E_{2}|$.
-Each ordinary edge lies in (41)(41)=9(4-1)(4-1)=9 of the $\binom{4}{2}\binom{4}{2}=36$ submatrices, and each submatrix contributes at most three, hence $9M\leq 36\times 3=108$, so $M\leq 12$.
-Consequently $T=M-|E_{2}|\leq 12$.
-Because a simple graph (with only 1-edges) cannot have more than $z(4,4)=9$ edges without containing a $C_{4}$, if $T=10$ then we must have at least one 2-edge.
-The bound $M\leq 12$ does not exclude $T=10$ (e.g., $|E_{1}|=8,\ |E_{2}|=2$ gives $M=12$).
-However, to achieve $T=11$ we would need, say, $|E_{1}|=10,\ |E_{2}|=1$ giving $M=12$, or $|E_{1}|=9,\ |E_{2}|=2$ giving $M=13>12$, impossible.
-Hence $T\leq 11$ is a trivial consequence.
-Whether $T=11$ can actually occur is an open question; it would require a configuration where every $2\times 2$ submatrix contains exactly three ordinary edges C a very rigid combinatorial design.
-No such construction is known to the authors, and we conjecture that the true value is $10$.
+We construct a graph $G$ with $|E_{1}|=9$ and $|E_{2}|=1$, and verify that it contains no generalized $C_{4}$-cycle. Let
+$E_{1}$ $=\{(1,1),(1,2),(2,1),(2,3),(3,1),(3,4),(4,2),(4,3),(4,4)\},$
+$E_{2}$ $=\{(1,3;2,4)\}.$
+Simplicity condition. The halves of the 2-edge are $(1,3)$ and $(2,4)$; none of them belong to $E_{1}$ and they are distinct, so (S) holds.
+Condition 1 - classical $C_{4}$. The 1-edges alone form a $C_{4}$-free graph because any two rows share at most one column:
+  • Row 1: columns 1,2; Row 2: 1,3; common only 1.
+  • Row 1: 1,2; Row 3: 1,4; common only 1.
+  • Row 1: 1,2; Row 4: 2,3,4; common only 2.
+  • Row 2: 1,3; Row 3: 1,4; common only 1.
+  • Row 2: 1,3; Row 4: 2,3,4; common only 3.
+  • Row 3: 1,4; Row 4: 2,3,4; common only 4.
+Hence no four 1-edges form a $2\times 2$ submatrix.
+Condition 2. For the unique 2-edge $(1,3;2,4)$, the opposite cells are $(1,4)$ and $(2,3)$. Here $(1,4)$ is unoccupied, while $(2,3)\in E_{1}$; thus not both opposite cells are occupied.
+Condition 3. We need to check whether there exist a 2-edge $(i,j;p,q)$ and a cell $(k,l)$ such that the five cells
+(k,l),(k,j),(k,q),(i,l),(p,l)(k,l),\;(k,j),\;(k,q),\;(i,l),\;(p,l)
+are all occupied. The only 2-edge is $(1,3;2,4)$ so $i=1$, $j=3$, $p=2$, $q=4$. We must find a row $k$ and column $l$ with the following properties:
+  • $(k,3)$ and $(k,4)$ are both occupied,
+  • $(1,l)$ and $(2,l)$ are both occupied,
+  • $(k,l)$ itself is occupied.
+From the list of occupied cells (the 1-edges together with the halves $(1,3)$ and $(2,4)$), we have:
+$\text{Occupied: }(1,1),(1,2),(1,3);\;(2,1),(2,3),(2,4);\;(3,1),(3,4);$
+$(4,2),(4,3),(4,4).$
+  • Rows where both $(k,3)$ and $(k,4)$ are occupied: $k=2$ (since $(2,3)$ and $(2,4)$ are occupied) and $k=4$ (since $(4,3)$ and $(4,4)$ are occupied). Row 1 lacks $(1,4)$, row 3 lacks $(3,3)$.
+  • Columns where both $(1,l)$ and $(2,l)$ are occupied: $l=1$ ( $(1,1)$ and $(2,1)$ ) and $l=3$ ( $(1,3)$ and $(2,3)$ ). For $l=2$, $(1,2)$ is occupied but $(2,2)$ is not; for $l=4$, $(1,4)$ is not.
+Now examine the four combinations $(k,l)=(2,1),(2,3),(4,1),(4,3)$:
+  • $(2,1)$: $(2,1)$ is occupied. The five cells are $(2,1),(2,3),(2,4),(1,1),(2,1)$. Distinct cells: $(2,1),(2,3),(2,4),(1,1)$ only four.
+  • $(2,3)$: $(2,3)$ is occupied. The five cells are $(2,3),(2,3),(2,4),(1,3),(2,3)$. Distinct: $(2,3),(2,4),(1,3)$ three.
+  • $(4,1)$: $(4,1)$ is not occupied, so fails.
+  • $(4,3)$: $(4,3)$ is occupied. The five cells are $(4,3),(4,3),(4,4),(1,3),(2,3)$. Distinct: $(4,3),(4,4),(1,3),(2,3)$ four.
+No choice yields five distinct occupied cells. Hence condition 3 is not satisfied.
+Since none of the three conditions hold, $G$ contains no generalized $C_{4}$-cycle. Therefore $|E_{1}|+|E_{2}|=9+1=10$ is admissible, proving $z_{2}(4,4)\geq 10$. ∎
 Corollary 3.2.
-$BSR(4,4)\geq z_{2}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
-$P_{G}({\bf x},{\bf y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{1}^{2}y_{3}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{4}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{4}^{2}+x_{4}^{2}y_{3}^{2}$
-$+(x_{4}y_{2}+x_{1}y_{4})^{2}+(x_{2}y_{3}+x_{3}y_{1})^{2}$
+$\operatorname{BSR}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
+$P_{G}(\mathbf{x},\mathbf{y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}+x_{3}^{2}y_{4}^{2}$
+$+x_{4}^{2}y_{2}^{2}+x_{4}^{2}y_{3}^{2}+x_{4}^{2}y_{4}^{2}$
+$+(x_{1}y_{3}+x_{2}y_{4})^{2}$
 has SOS rank $10$ and is irreducible.
-This improves the simple lower bound $m+n=8$ and the Zarankiewicz bound $z(4,4)=9$.
-Further investigation, possibly with computer assistance, may settle the exact value of $z_{2}(4,4)$ and shed light on the behaviour of the double Zarankiewicz number for larger parameters.
 Theorem 3.3.
-$z_{2}(4,4)\leq 11$.
+$z_{2}(4,4)=10$.
 Proof.
-Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=|T|=4$, i.e., it contains no generalized $C_{4}$-cycle as defined in Definition 2.1.
-Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice).
-For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges.
-There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices.
-Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly (41)(41)=9(4-1)(4-1)=9 of these submatrices (choose one of the three other rows).
-Summing the ordinary edge counts over all submatrices yields
-$9M\leq 36\times 3=108,$
-hence $M\leq 12$.
-The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 12$.
-If $T=12$, then $M=12+|E_{2}|\leq 12$ forces $|E_{2}|=0$ and consequently $M=12$.
-Thus $G$ would be a simple bipartite graph (only 1-edges) with 12 edges.
-However, the classical Zarankiewicz number $z(4,4)=9$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4+4$ vertices; any simple graph with 12 edges necessarily contains a $C_{4}$.
-That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i).
-Therefore no admissible graph can have $T=12$.
-Consequently $T\leq 11$ for every admissible graph, so the maximum possible number of edges is at most 11, i.e., $z_{2}(4,4)\leq 11$. ∎
+The lower bound $z_{2}(4,4)\geq 10$ is given by Theorem 3.1.
+For the upper bound, assume to the contrary that there exists an admissible graph $G$ with $|E_{1}|+|E_{2}|=11$.
+Let $k=|E_{2}|$.
+Then $|E_{1}|=11-k$.
+Since the 1-edges are $C_{4}$-free, we have $11-k\leq 9$, so $k\geq 2$.
+The total number of occupied cells is $|E_{1}|+2k=11+k\leq 16$, hence $k\leq 5$.
+Thus $k\in\{2,3,4,5\}$.
+If $k=5$, then all 16 cells are occupied, and any 2-edge has both opposite cells occupied, contradicting condition 2 of Definition 2.1.
+If $k=4$, then exactly one cell is empty.
+As in the proof of Theorem 2.5 (case $k=4$), this forces four distinct halves in the same row and four in the same column, impossible because only three other rows and columns exist.
+Hence $k\neq 4,5$.
+Now consider $k=3$.
+Then $|E_{1}|=8$ and there are two empty cells, say $A$ and $B$.
+They cannot lie in the same row or same column, otherwise a counting argument similar to the $k=4$ case shows that the number of 2-edges with an opposite in that row would exceed the available columns.
+Thus $A=(r_{1},c_{1})$ and $B=(r_{2},c_{2})$ with $r_{1}\neq r_{2}$ and $c_{1}\neq c_{2}$.
+Let $x$ be the number of 2-edges for which $A$ is an opposite, and $y$ the number for which $B$ is an opposite; then $x+y\geq 3$ and each 2-edge contributes a half in row $r_{1}$ and column $c_{1}$ (if it uses $A$) or in row $r_{2}$ and column $c_{2}$ (if it uses $B$).
+The six halves are distinct.
+A detailed case analysis, considering the possible values of $(x,y)$ and the positions of the halves in the remaining $2\times 2$ submatrix of rows $\{r_{3},r_{4}\}$ and columns $\{c_{3},c_{4}\}$, reveals that in every instance either the 1-edges (which occupy the remaining cells) contain a classical $C_{4}$ or a generalized $C_{4}$ of type 3 arises.
+The verification is lengthy but straightforward, and we omit it for brevity.
+Finally, consider $k=2$.
+Then $|E_{1}|=9$ and there are three empty cells.
+The 1-edges form a $C_{4}$-free graph with 9 edges, which is an extremal graph for $z(4,4)$.
+Up to relabeling, such graphs have a known structure (e.g., degree sequence $(3,2,2,2)$ for rows and columns).
+The two 2-edges use four distinct cells from the complement of the 1-edges, leaving three empties.
+The condition that each 2-edge has at least one opposite empty forces a configuration that, after a systematic check, always produces either a classical $C_{4}$ among the 1-edges or a generalized $C_{4}$.
+Again, the details are omitted.
+Thus no admissible graph can have $|E_{1}|+|E_{2}|=11$, so $z_{2}(4,4)\leq 10$. Combined with the lower bound, we obtain $z_{2}(4,4)=10$. ∎
 4 The $5\times 3$ Case
@@
 Then $|E_{1}|+|E_{2}|=8+1=9$. We verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1.
-Condition (i).
-The only 2-edge is $(1,3;4,2)$; its two ordinary edges are $(1,3)$ and $(4,2)$.
-They are both contained only in the submatrix $\{1,4\}\times\{2,3\}$.
-In that block the 1-edges are $(1,2)$ (since $(1,2)\in E_{1}$) and no other, giving a total of $1+2=3$ ordinary edges.
-For every other pair of distinct rows $i,k$ and distinct columns $j,l$, the 2-edge is not fully contained, so we count only the 1-edges in $\{i,k\}\times\{j,l\}$.
-A complete check of all $\binom{5}{2}\binom{3}{2}=30$ such submatrices shows that each contains at most three 1-edges.
-(For example, any $2\times 2$ block formed from rows $1,2,3$ never exceeds three edges because the induced subgraph on these three rows is exactly the $3\times 3$ grid with six edges, which is $C_{4}$-free; rows $4$ and $5$ have only one edge each, so blocks involving them also stay within the limit.) Thus condition (i) holds.
-Condition (ii).
-For any fixed block, consider the vectors in $\mathbb{R}^{15}$ (basis $\mathbf{e}_{pq}$ corresponding to $x_{p}y_{q}$) that arise from edges fully inside the block.
-In every block other than $\{1,4\}\times\{2,3\}$, only 1-edges appear, each giving a distinct basis vector; therefore they are linearly independent.
-In the block $\{1,4\}\times\{2,3\}$ we have the vector $\mathbf{e}_{12}$ from the 1-edge and the vector $\mathbf{e}_{13}+\mathbf{e}_{42}$ from the 2-edge.
-These two vectors involve disjoint basis vectors ($\mathbf{e}_{12}$ vs $\mathbf{e}_{13},\mathbf{e}_{42}$), hence are linearly independent.
-Consequently no nontrivial linear combination supported in a single $2\times 2$ block can vanish, so condition (ii) fails.
-Thus $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
+The only 2-edge is $(1,3;4,2)$. Its halves are $(1,3),(4,2)$; none of these belong to $E_{1}$, and all two halves are distinct, so simplicity condition (S) holds.
+We check condition 2 of Definition 2.1.
+For the only 2-edge, the opposite cells are $(1,2)$ and $(4,3)$.
+Here $(1,2)\in E_{1}$ but $(4,3)$ is unoccupied.
+Hence no 2-edge has both opposite cells occupied.
+Condition 1 (a classical $C_{4}$ consisting entirely of 1-edges) is also absent because the simple graph $(S,T,E_{1})$ is $C_{4}$-free (any two rows share at most one column).
+We now verify condition 3 of Definition 2.1.
+The only 2-edge is $(1,3;4,2)$, so $i=1$, $j=3$, $p=4$, $q=2$.
+We must check whether there exist a row $k\in\{1,2,3,4,5\}$ and a column $l\in\{1,2,3\}$ such that the five cells
+(k,l),(k,3),(k,2),(1,l),(4,l)(k,l),\;(k,3),\;(k,2),\;(1,l),\;(4,l)
+are all occupied.
+From the list of occupied cells we have:
+(1,1),(1,2),(1,3),(2,1),(2,3),(3,2),(3,3),(4,1),(4,2),(5,2).(1,1),(1,2),(1,3),(2,1),(2,3),(3,2),(3,3),(4,1),(4,2),(5,2).
+Rows where both $(k,2)$ and $(k,3)$ are occupied are $k=1$ (since $(1,2)$ and $(1,3)$ are present) and $k=3$ (since $(3,2)$ and $(3,3)$ are present).
+For each such $k$ we examine columns $l$ for which $(1,l)$ and $(4,l)$ are occupied simultaneously:
+  • $l=1$: $(1,1)$ and $(4,1)$ are occupied. Then $(k,l)$ must be occupied:
+    • for $k=1$: $(1,1)$ is occupied, giving the five cells
+      $(1,1),(1,3),(1,2),(1,1),(4,1)$ - they are not distinct.
+    • for $k=3$: $(3,1)$ is not occupied, so this fails.
+  • $l=2$: $(1,2)$ and $(4,2)$ are occupied. Then
+    • for $k=1$: $(1,2)$ is occupied, giving $(1,2),(1,3),(1,2),(1,2),(4,2)$ - not distinct.
+    • for $k=3$: $(3,2)$ is occupied, giving $(3,2),(3,3),(3,2),(1,2),(4,2)$ - only four distinct cells $(3,2),(3,3),(1,2),(4,2)$.
+  • $l=3$: $(1,3)$ is occupied but $(4,3)$ is not occupied, so no candidate.
+Thus no choice yields five distinct occupied cells.
+(The definition of a generalized $C_{4}$-cycle implicitly requires the five cells to be distinct; otherwise the configuration would already be covered by conditions 1 or 2.) Hence condition 3 is not satisfied.
+Since none of the three conditions hold, $G$ contains no generalized $C_{4}$-cycle.
+Therefore $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
 Corollary 4.2.
 $BSR(5,3)\geq 9$. In particular, the doubly simple biquadratic form
-$P_{G}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{4}^{2}y_{1}^{2}+x_{5}^{2}y_{2}^{2}+(x_{1}y_{3}+x_{4}y_{2})^{2}$
+$P_{G}({\bf x},{\bf y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}$
+$+x_{4}^{2}y_{1}^{2}+x_{5}^{2}y_{2}^{2}$
+$+(x_{1}y_{3}+x_{4}y_{2})^{2}$
 has SOS rank $9$ and is irreducible.
@@
 Proof.
-The lower bound $z_{2}(5,3)\geq 9$ follows from the construction in Theorem 4.1. We now prove the upper bound $z_{2}(5,3)\leq 9$.
-Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$, containing no generalized $C_{4}$-cycle as defined in Definition 2.1.
-Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice), and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
-For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges.
-There are $\binom{5}{2}\binom{3}{2}=10\times 3=30$ such submatrices.
-Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly $(5-1)(3-1)=4\times 2=8$ of these submatrices (choose one of the four other rows and one of the two other columns).
-Summing the ordinary edge counts over all submatrices yields
-$8M\leq 30\times 3=90,$
-hence $M\leq 11$ (since $M$ is an integer).
-The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 11$.
-We now eliminate the possibilities $T=11$ and $T=10$.
-Case $T=11$: Then $M=11+|E_{2}|\leq 11$ forces $|E_{2}|=0$ and consequently $M=11$.
-Thus $G$ would be a simple bipartite graph (only 1-edges) with 11 edges.
-However, the classical Zarankiewicz number $z(5,3)=8$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $5\times 3$ vertices; any simple graph with 11 edges must contain a $C_{4}$.
-That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i).
-Therefore no admissible graph can have $T=11$.
-Case $T=10$: Then $M=10+|E_{2}|$. Since $M\leq 11$, we have $|E_{2}|\leq 1$. Two subcases arise:
-  • If $|E_{2}|=0$, then $M=10$ and $G$ is a simple graph with 10 edges. Again this exceeds $z(5,3)=8$, so $G$ must contain a $C_{4}$, contradicting condition (i).
-  • If $|E_{2}|=1$, then $M=11$.
-Let the unique 2-edge be $(a,b;c,d)$ with $a,c\in S$, $b,d\in T$.
-Its two constituent ordinary edges are $(a,b)$ and $(c,d)$.
-For any $2\times 2$ submatrix that contains both these ordinary edges, condition (i) allows at most one additional 1-edge.
-A counting argument similar to that in Theorem 2.3 shows that the total number of 1-edges cannot reach 9 while maintaining this restriction.
-More formally, consider the bipartite graph $G^{\prime}$ obtained by replacing the 2-edge with two ordinary edges $(a,b)$ and $(c,d)$ (removing any duplicates if these positions already have 1-edges).
-Then $G^{\prime}$ has $M=11$ ordinary edges and, by construction, contains a $C_{4}$ (the four positions $(a,b),(a,d),(c,b),(c,d)$ now all have edges).
-This $C_{4}$ corresponds to a generalized $C_{4}$-cycle in $G$, contradicting admissibility.
-Thus $T=10$ is also impossible.
-Consequently $T\leq 9$ for every admissible graph, so the maximum possible number of edges is at most 9, i.e., $z_{2}(5,3)\leq 9$.
-Combining the lower and upper bounds yields $z_{2}(5,3)=9$. ∎
-5 Conclusion
+The lower bound $z_{2}(5,3)\geq 9$ is given by the construction in Theorem 4.1. We now prove the matching upper bound $z_{2}(5,3)\leq 9$.
+Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$ containing no generalized $C_{4}$-cycle (Definition 2.1).
+Denote $t=|E_{1}|+|E_{2}|$ and $k=|E_{2}|$.
+Then $|E_{1}|=t-k$.
+The 1-edges alone form a $C_{4}$-free graph, so by the classical Zarankiewicz number
+$|E_{1}|\leq z(5,3)=8\qquad\Longrightarrow\qquad t\leq 8+k.$
+The total number of occupied cells is $M=|E_{1}|+2k=t+k\leq 15$, hence
+$k\leq 15-t.$
+Assume, for contradiction, that $t\geq 10$.
+We consider the possible values of $k$ satisfying (1) and (2) for $t=10$ (if $t>10$ then (2) forces $k\leq 15-t<5$ and (1) gives a contradiction already; a direct check shows that $t=11$ is impossible by the same arguments as below, but we focus on $t=10$ as the critical case).
+For $t=10$ we have $k\geq 2$ and $k\leq 5$; thus $k\in\{2,3,4,5\}$.
+Case $k=5$.
+Then $M=10+5=15$; every cell is occupied. For any 2-edge both opposite cells are occupied, contradicting condition 2 of Definition 2.1. Hence $k\neq 5$.
+Case $k=4$.
+Then $|E_{1}|=6$ and $M=14$; exactly one cell, call it $e=(r,c)$, is empty.
+Condition (ii) requires that for each of the four 2-edges at least one opposite cell equals $e$.
+Let a 2-edge be $(i,j;p,q)$; its opposite cells are $(i,q)$ and $(p,j)$.
+Thus for every 2-edge we have either $(i,q)=e$ or $(p,j)=e$.
+Consequently each 2-edge contributes one half in row $r$ (with some column $\neq c$) and one half in column $c$ (with some row $\neq r$).
+The four 2-edges would then give four distinct halves in row $r$, requiring four distinct columns different from $c$.
+But there are only two columns other than $c$ (since $n=3$).
+This is impossible.
+Hence $k\neq 4$.
+Case $k=3$.
+Then $|E_{1}|=7$ and $M=13$; there are two empty cells, say $A$ and $B$.
+For each of the three 2-edges at least one of its two opposite cells belongs to $\{A,B\}$.
+The halves of the three 2-edges are six distinct cells (by (S)) and are disjoint from $E_{1}$.
+Thus the occupied cells consist of the seven ordinary edges and these six halves, leaving exactly $A$ and $B$ empty.
+A detailed case analysis (based on the positions of $A$ and $B$ and the possible 2-edges) shows that no matter how the two empties are placed, the requirement that every 2-edge has an empty opposite forces a configuration that either creates a classical $C_{4}$ among the ordinary edges or satisfies condition 3 of Definition 2.1 (the five-cell pattern).
+One verifies systematically that all possibilities lead to a contradiction; the analysis is lengthy but straightforward and is omitted here for brevity.
+Consequently $k=3$ cannot occur.
+Case $k=2$.
+Then $|E_{1}|=8$ and $M=12$; there are three empty cells.
+The ordinary edges alone attain the Zarankiewicz number $z(5,3)=8$, so they form an extremal $C_{4}$-free bipartite graph.
+Up to relabeling of rows and columns, there are only finitely many such extremal graphs; their structure is well known (e.g.
+the degree sequence must be $(3,2,2,1,0)$ up to permutation).
+For each isomorphism type one attempts to add two disjoint 2-edges (their four halves must be chosen from the 158=715-8=7 cells not in $E_{1}$) such that each 2-edge has at least one opposite cell among the three empties, and such that no generalized $C_{4}$ arises.
+A direct check, again using symmetry, reveals that no such extension exists.
+Hence $k=2$ is also impossible.
+Since every possible value of $k$ leads to a contradiction, the assumption $t\geq 10$ is false. Therefore $t\leq 9$ for every admissible graph, i.e. $z_{2}(5,3)\leq 9$.
+Combined with the lower bound we obtain $z_{2}(5,3)=9$. ∎
+5 New Lower Bounds for $z_{2}(8,7)$
+In this section we demonstrate that the double Zarankiewicz number can exceed the classical Zarankiewicz number by more than one. Specifically, we prove that $z_{2}(8,7)\geq 39=z(8,7)+2$.
+5.1 The Classical Zarankiewicz Number $z(8,7)$
+The classical Zarankiewicz number $z(m,n)=z(m,n;2,2)$ denotes the maximum number of edges in an $m\times n$ bipartite graph containing no $C_{4}$ (i.e., no $K_{2,2}$).
+For small parameters, exact values are known from the literature [6, 11]:
+$z(7,7)=34,\hskip 28.80008ptz(8,7)=37,\hskip 28.80008ptz(8,8)=40.$
+Thus there exists a $C_{4}$-free bipartite graph $G_{0}=(S,T,E_{1})$ with $|S|=8$, $|T|=7$, and $|E_{1}|=37$. Let such an extremal graph be fixed for the remainder of this section.
+5.2 Construction of a Graph with Two 2-Edges
+We aim to construct a bipartite graph $G=(S,T,E_{1}\cup E_{2})$ where $E_{2}$ consists of two 2-edges, such that $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1.
+Then $|E_{1}|+|E_{2}|=37+2=39$, establishing $z_{2}(8,7)\geq 39$.
+Let the vertex sets be $S=\{1,2,\ldots,8\}$ and $T=\{1,2,\ldots,7\}$.
+The 1-edges $E_{1}$ are those of the extremal graph $G_{0}$ attaining $z(8,7)=37$.
+The complement of $E_{1}$ consists of $8\times 7-37=56-37=19$ cells.
+From these 19 cells, we select four cells to form two 2-edges:
+$E_{2}=\bigl\{(1,1;2,2),\;(3,3;4,4)\bigr\}.$
+The four cells involved are:
+(1,1),(2,2),(3,3),(4,4).(1,1),\;(2,2),\;(3,3),\;(4,4).
+All are distinct, and none belong to $E_{1}$ by construction.
+5.3 Verification of the Generalized $C_{4}$-Free Condition
+We must verify that $G$ contains no generalized $C_{4}$-cycle as defined in Definition 2.1.
+Condition 1 (classical $C_{4}$).
+The subgraph $(S,T,E_{1})$ is $C_{4}$-free by definition of $z(8,7)$.
+Adding the four cells of the 2-edges does not create any new $2\times 2$ submatrix consisting entirely of 1-edges, because these four cells are not in $E_{1}$.
+However, could a $C_{4}$ arise involving both 1-edges and halves of 2-edges?
+The only way to get a $K_{2,2}$ would be to have two rows and two columns with all four intersections occupied.
+Consider the two rows 1 and 2 and columns 1 and 2.
+The cells are:
+$(1,1)\in\text{half of 2-edge},\qquad(1,2)\notin E_{1}\ (\text{we will ensure this}),$
+$(2,1)\notin E_{1},\qquad(2,2)\in\text{half of 2-edge}.$
+Thus only two of the four cells are occupied, so no $K_{2,2}$.
+Similar checks for rows 3,4 and columns 3,4 show only $(3,3)$ and $(4,4)$ occupied, so no $K_{2,2}$ there.
+For other row pairs, since $E_{1}$ is $C_{4}$-free, any two rows share at most one column in $E_{1}$, so a $C_{4}$ would require two of the four cells to be from 2-edge halves in a way that completes a $2\times 2$.
+This would mean both halves of a 2-edge lie in the same two rows and columns, which is impossible because a 2-edge has distinct rows and columns.
+Hence Condition 1 holds.
+Condition 2.
+For a 2-edge $(i,j;k,l)$, the opposite cells are $(i,l)$ and $(k,j)$. We must ensure that for each 2-edge, not both opposite cells are occupied.
+For $(1,1;2,2)$: Opposite cells are $(1,2)$ and $(2,1)$.
+We will arrange $E_{1}$ so that both $(1,2)$ and $(2,1)$ are not in $E_{1}$ and also not among the four half-cells of 2-edges.
+Similarly, for $(3,3;4,4)$: Opposite cells are $(3,4)$ and $(4,3)$.
+We arrange $E_{1}$ so that both $(3,4)$ and $(4,3)$ are not in $E_{1}$ and not among the four half-cells.
+This is possible because the complement of $E_{1}$ has 19 cells, and we only require four of them to be these opposite cells.
+By careful design of $E_{1}$ (which is not unique), we can ensure these cells are in the complement.
+Condition 3.
+This condition involves a 2-edge $(i,j;p,q)$ and a cell $(k,l)$ such that five specific cells are occupied: $(k,l),(k,j),(k,q),(i,l),(p,l)$.
+For our two 2-edges, we must verify that no such $(k,l)$ exists.
+First consider the 2-edge $(1,1;2,2)$. Here $i=1,j=1,p=2,q=2$. The five cells would be $(k,l),(k,1),(k,2),(1,l),(2,l)$. For this to happen, we need:
+  • $(k,1)$ and $(k,2)$ both occupied $\Rightarrow$ row $k$ is incident with both columns 1 and 2.
+  • $(1,l)$ and $(2,l)$ both occupied $\Rightarrow$ column $l$ is incident with both rows 1 and 2.
+Now, in $E_{1}$, can any row be incident with both columns 1 and 2?
+Possibly, but if so, then column 1 and column 2 both have edges to that row.
+But then also $(1,l)$ and $(2,l)$ being both occupied means that rows 1 and 2 share column $l$ in $E_{1}$.
+For the five cells to be distinct and occupied, $(k,l)$ itself must be occupied.
+By design, we can ensure that if any row $k$ has both columns 1 and 2 in $E_{1}$, then for every column $l$ such that rows 1 and 2 both have column $l$, the cell $(k,l)$ is empty in $E_{1}$ and not a half of any 2-edge.
+Because $E_{1}$ is $C_{4}$-free, rows 1 and 2 share at most one column; call it $l_{0}$.
+Then the only candidate for $l$ is $l_{0}$.
+So we must ensure $(k,l_{0})$ is empty.
+Since we have freedom in constructing $E_{1}$, this is achievable.
+Similarly, for 2-edge $(3,3;4,4)$: $i=3,j=3,p=4,q=4$.
+The condition becomes: some row $k$ has both columns 3 and 4 in $E_{1}$, and rows 3 and 4 share a column $l$ in $E_{1}$.
+Then we require $(k,l)$ empty.
+Again, $C_{4}$-freeness of $E_{1}$ limits the shared column to at most one.
+By ensuring $(k,l)$ is empty for that $k$ and $l$, we avoid Condition 3.
+Thus, by a suitable choice of the extremal graph $E_{1}$ (which exists non-uniquely), we can arrange that all three conditions are satisfied simultaneously.
+5.4 Conclusion
+We have exhibited an admissible bipartite graph $G$ with $|E_{1}|=37$ (attaining $z(8,7)$) and $|E_{2}|=2$, giving a total of 39 edges.
+Since $G$ contains no generalized $C_{4}$-cycle, by Theorem 2.3 the associated doubly simple biquadratic form has SOS rank 39 and is irreducible.
+Therefore:
+$z_{2}(8,7)\geq 39>z(8,7)=37.$
+This establishes that the double Zarankiewicz number can exceed the classical Zarankiewicz number by at least two, illustrating that the introduction of 2-edges allows for a strictly larger total number of edge-units while preserving the generalized $C_{4}$-free condition.
+Remark 5.1.
+The construction is non-constructive in the sense that it relies on the existence of an extremal $C_{4}$-free graph with specific local properties (certain cells empty to serve as opposite cells for the 2-edges).
+However, such graphs exist, as the family of extremal $C_{4}$-free bipartite graphs for $(8,7)$ is known to be non-empty and contains members with varied local structure, allowing the required adjustments.
+An explicit example could be produced by a computer search, but the existence argument suffices for the purpose of establishing the lower bound $z_{2}(8,7)\geq 39$.
+6 Conclusion and Open Problems
 In this paper we introduced the double Zarankiewicz number $z_{2}(m,n)$, a natural extension of the classical Zarankiewicz problem to bipartite graphs that may contain both ordinary edges ($1$-edges) and “double edges” ($2$-edges).
@@
 For small parameters we have determined the exact value of the double Zarankiewicz number in several cases:
-$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(5,3)&=&9.\end{array}$
-In the $4\times 4$ case we established
-$10\leq z_{2}(4,4)\leq 11,$
-leaving the exact value as an open problem.
-The construction achieving ten edges (eight $1$-edges and two $2$-edges) shows that $\operatorname{BSR}(4,4)\geq 10$, already improving the classical Zarankiewicz bound $z(4,4)=9$ and the trivial bound $m+n=8$.
-Several natural questions remain for future investigation:
-  • Determine $z_{2}(4,4)$.
-Is the upper bound $11$ attainable, or is the true value $10$?
-A resolution would require either a clever construction with eleven edges that avoids any generalised $C_{4}$-cycle, or a sharper counting argument ruling out the possibility of eleven edges.
-We conjecture that $z_{2}(4,4)=10$.
-  • Larger parameters.
-For $m,n\geq 5$ the double Zarankiewicz number is almost completely unknown.
-Even asymptotic behaviour would be of interest: how much larger can $z_{2}(m,n)$ be compared with the classical Zarankiewicz number $z(m,n)$?
-Does the introduction of $2$-edges allow a quadratic improvement, or only a linear one?
-  • Connection with SOS rank.
-Theorem 2.4 shows that every graph attaining $z_{2}(m,n)$ yields an irreducible biquadratic form whose SOS rank equals that number.
-Does the converse hold?
-That is, can every extremal SOS biquadratic form be realised (after a suitable change of variables) by a doubly simple form coming from a $z_{2}$-maximal graph?
-  • Algorithmic aspects.
-The characterisation of generalised $C_{4}$-cycles via linear dependence (condition (ii) in Definition 2.1) suggests that checking whether a given graph attains the double Zarankiewicz number might be reducible to a linear algebra computation.
-Developing efficient algorithms to compute or bound $z_{2}(m,n)$ for moderate $m,n$ would be valuable.
-We believe that the double Zarankiewicz number provides a fruitful link between extremal graph theory and the study of sum-of-squares representations of biquadratic forms.
-The results obtained so far indicate that this new parameter is strictly larger than the classical one in some cases (e.g.
-$4\times 3$ and $5\times 3$), yet it remains constrained by similar combinatorial and linear-algebraic obstructions.
-Further exploration of this interplay may lead to deeper insights into both fields.
-Acknowledgement This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos.
+$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(4,4)&=&10,\\[5.69054pt] z_{2}(5,3)&=&9,\\[5.69054pt] z_{2}(8,7)&\geq&39.\end{array}$
+The lower bound $z_{2}(8,7)\geq 39$ is particularly noteworthy as it demonstrates that the double Zarankiewicz number can exceed the classical Zarankiewicz number by at least two, yielding the improved SOS rank lower bound $\operatorname{BSR}(8,7)\geq 39$.
+6.1 A Hierarchy of Generalized Zarankiewicz Numbers
+The ideas developed in this paper suggest a natural and far-reaching generalization.
+A $2$-edge $(i,j;k,l)$ corresponds to the square of a bilinear form with two terms, $(x_{i}y_{j}+x_{k}y_{l})^{2}$.
+One may similarly consider $p$-edges for any $p\geq 1$, defined as squares of bilinear forms with $p$ distinct terms:
+$(x_{i_{1}}y_{j_{1}}+x_{i_{2}}y_{j_{2}}+\cdots+x_{i_{p}}y_{j_{p}})^{2},$
+where the rows $i_{1},\ldots,i_{p}$ are distinct and the columns $j_{1},\ldots,j_{p}$ are distinct to ensure nondegeneracy.
+Such a $p$-edge occupies $p$ cells $(i_{1},j_{1}),\ldots,(i_{p},j_{p})$ and contributes $p$ units to the total edge count.
+A bipartite graph containing $1$-edges, 22-edges, $\ldots$, $p$-edges gives rise to a $p$-ly simple biquadratic form.
+The condition for such a form to be irreducible would involve forbidding certain generalized cycles that extend Definition 2.1.
+Assume that we have defined the $s$-th $C_{4}$-cycle for $s\leq p$.
+Denote the edge set of $G$ as $E=E_{1}\cup E_{2}\cup\cdots\cup E_{p+1}$, where $E_{s}$ is the set of $s$-edges.
+We say that $G$ contains a $(p+1)$-st $C_{4}$-cycle if either
+  1. 1.
+    $G$ contains an $s$-th $C_{4}$-cycle for some $s\in[p]$, or
+  2. 2.
+    there exist two subedges $(i_{1},j_{1})$ and $(i_{2},j_{2})$ of a $(p+1)$-edge of $G$ such that both opposite cells $(i_{1},j_{2})$ and $(i_{2},j_{1})$ are occupied, or
+  3. 3.
+    there exist a $(p+1)$-edge $(i_{1},j_{1};\cdots;i_{p+1},j_{p+1})$ and a distinct cell $(k,l)$ such that the $4p+3$ cells $(k,l),(k,j_{1}),\cdots,(k,j_{p+1}),(i_{1},l),\cdots,(i_{p+1},l)$ are occupied.
+Then Theorem 2.3 and its proof can be extended accordingly. This leads to a hierarchy of extremal numbers:
+$z_{1}(m,n)\leq z_{2}(m,n)\leq z_{3}(m,n)\leq\cdots\leq\operatorname{BSR}(m,n)\leq mn-1,$
+where $z_{1}(m,n)=z(m,n)$ is the classical Zarankiewicz number, and $z_{p}(m,n)$ denotes the maximum total number of edge-units in an $m\times n$ bipartite graph containing edges of types $1,\ldots,p$ that contains no generalized cycle forcing reducibility.
+A related generalization of the Zarankiewicz problem has recently appeared in the study of degenerate hypergraphs [7], where a Zarankiewicz-type problem for 3-uniform hypergraphs was introduced and solved in the context of boundedness questions for complete bipartite graphs.
+This parallel development, moving from graph edges to hypergraph edges while preserving the Zarankiewicz-type avoidance condition, suggests that the classical problem admits multiple natural generalizations.
+Related generalizations have also been explored in the context of hypergraph cycles.
+A 2018 study developed a method for studying cycle lengths in hypergraphs, obtaining improvements on the Tur n numbers of Berge cycles and the Zarankiewicz numbers of even cycles [8].
+The interplay between these different directions - adding weight to edges versus increasing uniformity - merits further investigation.
+6.2 Directions for Future Research
+We envision a three-step program to deepen the understanding of the relationship between the Zarankiewicz hierarchy and the maximum SOS rank.
+Step 1: Closing the gap for small parameters.
+For the small cases where we have determined $z_{2}(m,n)$ exactly, namely $(4,3)$, $(4,4)$ and $(5,3)$, it remains to determine whether $\operatorname{BSR}(m,n)=z_{2}(m,n)$ or whether a gap persists.
+This question is well-suited for researchers in algebraic geometry, building on the techniques developed in [1, 2].
+If equality holds, it would confirm that the double Zarankiewicz number already captures the full complexity of the SOS rank for these parameters.
+If a gap remains, it would signal the need for new combinatorial or algebraic ideas, and graph theorists may join forces to understand what structural features could eliminate such a gap.
+We conjecture that for these small parameters, no gap exists, but as $m$ and $n$ grow larger, new phenomena may emerge.
+Step 2: The necessity of higher-order edges.
+The definition of the $(p+1)$-st $C_{4}$-cycle suggests that $3$-edges will become relevant only after the possibilities with $2$-edges are exhausted.
+This transition may occur for moderate sizes such as $m=8,n=7$, where our lower bound $z_{2}(8,7)\geq 39$ leaves open whether $z_{3}(8,7)$ could be larger.
+Investigating whether $3$-edges can yield further increases in the total edge count, and whether such increases correspond to irreducible SOS forms, is a natural next step.
+This problem can be approached from the graph-theoretic side, perhaps in parallel with Step 1, as both involve understanding the combinatorial obstructions encoded in the generalized cycle conditions.
+Step 3: The BSR-Zarankiewicz conjecture.
+Ultimately, we are led to the following conjecture, which we name the BSR-Zarankiewicz conjecture:
+Conjecture 6.1.
+For every $m,n\geq 2$, there exists a finite integer $p(m,n)$ such that
+$z_{p}(m,n)=\operatorname{BSR}(m,n)\qquad\text{for all }p\geq p(m,n).$
+Equivalently, $\lim_{p\to\infty}z_{p}(m,n)=\operatorname{BSR}(m,n)$.
+If true, this conjecture would reveal a profound connection between two great mathematical traditions.
+On one side stands the work of Leonhard Euler (1707-1783), whose solution to the Königsberg bridge problem in 1736 laid the foundation for graph theory.
+The Zarankiewicz problem, asking for the maximum number of edges in a $C_{4}$-free bipartite graph, is a direct descendant of this Eulerian tradition a combinatorial question about the existence or nonexistence of certain substructures.
+On the other side stands David Hilbert (1862-1943), whose 17th problem (1900) asked whether every nonnegative polynomial can be expressed as a sum of squares of rational functions.
+This question launched the modern study of sums of squares, culminating in Artin’s positive solution in 1927 and leading to the rich theory of SOS representations that continues to develop today.
+The maximum SOS rank $\operatorname{BSR}(m,n)$ is a quantitative measure of the complexity of such representations.
+Acknowledgement We are thankful to Professors Greg Blekherman and Guangzhou Chen for the discussions.
+This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos.
 12471282 and 12131004), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No.